### v1

# Synergizing Foundation Models and Federated Learning: A Survey

Abstract
The recent development of Foundation Models (FMs), represented by large language models, vision transformers, and multimodal models, has been making a significant impact on both academia and industry. Compared with small-scale models, FMs have a much stronger demand for high-volume data during the pre-training phase. Although general FMs can be pre-trained on data collected from open sources such as the Internet, domain-specific FMs need proprietary data, posing a practical challenge regarding the amount of data available due to privacy concerns. Federated Learning (FL) is a collaborative learning paradigm that breaks the barrier of data availability from different participants. Therefore, it provides a promising solution to customize and adapt FMs to a wide range of domain-specific tasks using distributed datasets whilst preserving privacy. This survey paper discusses the potentials and challenges of synergizing FL and FMs and summarizes core techniques, future directions, and applications. A periodically updated paper collection on FM-FL is available at https://github.com/lishenghui/awesome-fm-fl.
1 Introduction
The landscape of Artificial Intelligence (AI) has been revolutionized by the emergence of Foundation Models (FMs) (Bommasani et al., 2021), such as BERT Devlin et al. (2019), GPT series Brown et al. (2020); OpenAI (2022, 2024), and LLaMA series Touvron et al. (2023a, b) in Natural Language Processing (NLP); ViTs Dosovitskiy et al. (2021) and SAM Kirillov et al. (2023) in Computer Vision (CV); CLIP Radford et al. (2021), DALL-E Ramesh et al. (2021), Gemini Google (2023), and GPT-4o in multimodal applications. These FMs have become pivotal in a myriad of AI applications across diverse domains. Their superb capability to generalize across tasks and domains stems from their pre-training on extensive datasets (Gunasekar et al., 2023), which imbues them with a profound understanding of language, vision, and multimodal data.
While general-purpose FMs can leverage openly accessible data from the Internet, domain-specific FMs require proprietary data. It is, however, challenging to collect vast amounts of proprietary data and perform centralized pre-training or fine-tuning for domain-specific FMs, due to privacy restrictions Jo and Gebru (2020); GDPR (2016); CCPA (2023). Particularly in domains such as law, healthcare, and finance, where data is inherently privacy-sensitive, there is a pressing need for stringent privacy safeguards. Furthermore, given that data often constitutes a pivotal asset for enterprises, its widespread distribution is prohibitive. Consequently, there is an urgent need for novel strategies to handle data availability and facilitate model training, thereby unlocking the potential of domain-specific FMs whilst respecting data privacy.
To address the challenges associated with data privacy in model training, Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising paradigm. FL facilitates collaborative model training across decentralized clients without the need to share raw data, thus ensuring privacy preservation. Concretely, FL encompasses periodic interactions between the server and decentralized clients for the exchange of trainable model parameters without the requirement for private client data. Recognizing such a benefit, integrating FMs with FL presents a compelling solution for domain-specific FMs Zhuang et al. (2023); Yu et al. (2023d).
Despite the potential synergies between FL and FMs, the field is still nascent, lacking a comprehensive understanding of challenges, methodologies, and directions. This survey aims to bridge this gap by providing a thorough exploration of the integration of FMs and FL. We delve into the motivations and challenges of combining these two paradigms, highlight representative techniques, and discuss applications and future directions. By elucidating the intersection of FL and FMs, we aim to catalyze further research and innovation in this burgeoning area, ultimately advancing the development of privacy-aware, domain-specific FMs.
The paper continues as follows: The next section introduces background on FMs and FL. Section 3 presents the motivation and challenges for synergizing FMs and FL. Section 4 highlights representative techniques. Section 5 explores the applications across various domains. Before concluding, we discuss representative future directions in Section 6.
2 Background
2.1 Foundation Models
An FM is a model that can be adapted to a wide array of tasks through fine-tuning after initial pre-training Bommasani et al. (2021). The lifecycle of FMs typically involves pre-training on extensive generic data to establish the basis of their abilities Bubeck et al. (2023), followed by adaptation to downstream tasks such as domain-specific question answering Zhang et al. (2023e), and ultimately application in various domains.
FMs have sparked a significant paradigm shift in various fields of AI such as NLP, CV, speech and acoustics, and beyond. In the realm of NLP, the most prominent example is Large Language Models (LLMs) with substantial parameter sizes (Zhao et al., 2023). These models, such as ChatGPT and GPT-4 (OpenAI, 2022, 2024), demonstrate exceptional abilities in natural language understanding and generation, enabling them to comprehend and respond to user inputs with remarkable contextual relevance. This capability proves invaluable in applications like customer service, virtual assistants, and chatbots, where effective communication is paramount. Moreover, LLMs eliminate the need for training models from scratch for specific tasks, be it machine translation, document summarization, text generation, or other language-related tasks.
[…]
2.2 Federated Learning
FL McMahan et al. (2017) is a learning paradigm that enables a collection of clients to collaboratively learn a shared global model by leveraging their private datasets in a distributed manner, assisted by the coordination of a central server. The general goal of FL is to find a parameter set $\bm{\theta}$ that minimizes the following distributed optimization objective:
$\min\limits_{\bm{\theta}}F(\bm{\theta}):=\frac{1}{K}\sum_{k\in[K]}F_{k}(\bm{\theta}),$ (1)
where $K$ represents the total number of clients and $F_{k}(\bm{\theta})=\mathbb{E}_{\bm{z}\sim\mathcal{D}_{k}}[\ell(\bm{\theta};\bm{z})]$ denotes the expected risk of the $k$-th client. Here, $\mathcal{D}_{k}$ is the data distribution for the $k$-th client, and $\ell(\cdot;\cdot)$ is a user-specified loss function.
The most representative algorithms in the FL literature are the FedAvg-family algorithms McMahan et al. (2017); Reddi et al. (2021). The standard FedAvg involves periodic interactions between the server and decentralized clients to exchange trainable model parameters. In this process, each client independently trains the model on its local data and sends the model updates to a central server. The server aggregates these updates by computing their average to update the global model, which is subsequently redistributed to the clients for further iterations. Many variants have been proposed to tackle issues such as convergence and local data heterogeneity Diao et al. (2021). For example, FedProx Li et al. (2020) and FedDyn Acar et al. (2021) introduce regularizer terms to penalize client updates that are far away from the server model. A general framework FedOpt Reddi et al. (2021) unifies adaptive optimizers (Adam, Yogi, etc.) and demonstrates superior convergence speed when compared to the naive FedAvg.
FL offers an efficient privacy-preserving way to train models on large-scale and diverse data Kairouz et al. (2021), leading to its application across various domains such as healthcare Lincy and Kowshalya (2020); Rieke et al. (2020); Joshi et al. (2022), finance Chatterjee et al. (2023); Liu et al. (2023b), and smart cities Ramu et al. (2022); Pandya et al. (2023).
3 FM-FL: Motivation & Challenges
In this section, we first motivate the synergy of FMs and FL (Section 3.1), then summarize the key challenges (Section 3.2).
3.1 Motivation
The integration of FMs and FL represents a compelling collaboration that leverages each other’s strengths to address their respective limitations, embodying a complementary relationship Zhuang et al. (2023); Li and Wang (2024).
FL expands data availability for FMs
By leveraging data from a wide range of sources in a privacy-preserving manner, FL makes it possible to build models on sensitive data in specific domains, such as healthcare Lincy and Kowshalya (2020); Joshi et al. (2022); Rieke et al. (2020) and finance Chatterjee et al. (2023); Liu et al. (2023b). This enhances the diversity and volume of training data, improving model robustness and adaptability. Moreover, FL enables the integration of personal and task-specific data, allowing FMs to be customized for personal applications. For instance, Google has trained next-word-prediction language models on mobile keyboard input data with FL to improve user experience Xu et al. (2023); Bonawitz et al. (2021).
FMs boost FL with feature representation and few-shot learning capabilities
By pre-training on large-scale generic data, FMs acquire essential knowledge and understanding capabilities Brown et al. (2020), providing multiple benefits to FL. Firstly, they benefit FL systems by offering advanced feature representations and learning capabilities from the outset. Secondly, leveraging the pre-learned knowledge of FMs can accelerate the FL process, enabling efficient and effective adaptation to specific tasks with minimal additional training. Thirdly, FMs’ powerful generative capabilities could help FL overcome the data heterogeneity challenge by synthesizing extra data, thus accelerating model convergence Huang et al. (2024).
3.2 Core Challenges
In this part, we discuss challenges emerging from the FM-FL marriage in three aspects: efficiency, adaptability, as well as trustworthiness.
Efficiency Challenges
Efficiency challenges stem from the mismatch between the significant resource demands of FM training and the limited, heterogeneous system resources (e.g., mobile devices) within FL systems, such as communication bandwidth, computational power, and memory Su et al. (2023). The communication bottleneck of FL is induced by frequently exchanging training information between the server and clients over limited bandwidth channels Kairouz et al. (2021). The substantial number of parameters in FMs further exacerbates this burden, thus hindering the training process.
Adaptability Challenges
Adaptability challenges arise from the adaptation of an FM to a specific downstream task (e.g., by fine-tuning) in FL settings. Key challenges include data heterogeneity and resource heterogeneity. Performance degradation in FL, attributed to heterogeneous data distributions among clients, is a well-recognized issue Kairouz et al. (2021); Li et al. (2022). A recent study Babakniya et al. (2023a) has shown that such performance penalty is even more substantial when fine-tuning FMs. For NLP tasks, data heterogeneity can manifest as variations in language, style, topic, or sentiment across datasets held by different clients. In multi-modal scenarios, the challenge is even more pronounced due to the inherent diversity in data types (e.g., text, images, and audio) Yu et al. (2023a). Addressing data heterogeneity involves not just identifying and measuring it but also developing algorithms that are robust to such diversity, ensuring that the model can learn effectively from varied data contributions without compromising on performance. In terms of resource heterogeneity, the memory and computational resources of the devices for different participants may be diverse Diao et al. (2021), which could cause delays for model synchronization and inactivation of some participants, i.e., stragglers, making it challenging to leverage the full potential of FMs in FL settings.
Trustworthiness Challenges
Trustworthiness challenges emphasize the concerns regarding privacy, security, and ethical considerations in the lifecycle of FM-FL, from the pre-training and model adaptation to the application stages. We present two representative challenges from this perspective: (1) intellectual property: Intellectual Property (IP) protection in FM-FL primarily involves attributing ownership rights for both models and data. From the server’s perspective, broadcasting a pre-trained model to multiple nodes for fine-tuning poses IP protection and security risks (e.g., model theft), necessitating measures to safeguard IP rights and ensure model integrity Kang et al. (2024); (2) privacy leakage: Although FL does not immediately share data, studies have shown that it may not always guarantee sufficient privacy preservation Geiping et al. (2020), as model parameters (e.g., weights or gradients) may leak sensitive information to malicious adversaries Zhu et al. (2019). (3)Poisoning Attacks: FL systems are inherently vulnerable to attacks due to their wide attack surface and reliance on network communication Li et al. (2023b). Poisoning attacks are carried out by malicious participants, aiming to bias the global model to the desire of attackers.
{forest}
[…]
4.1), l1node, [Parameter-Efficient
Fine-Tuning, l2node, [Selective, l3node, [  RaFFM Yu et al. (2023c), FedBF Zhang et al. (2023f) , leaf] ] [Additive, l3node, [  FedCLIP Lu et al. (2023a), FedDAT Chen et al. (2024) , leaf] ] [Reparameterization-based, l3node, [  HetLoRA Cho et al. (2024), FedDPA Yang et al. (2024b), leaf ] ] ] [Model Compression, l2node, [Sparsification, l3node, [ PruneFL Jiang et al. (2023c), FLASH Babakniya et al. (2023b),leaf] ] [Quantization, l3node, [ FedSplitBERT Lit et al. (2022),leaf] ] ] [Zeroth-Order
Optimization, l2node, [  BAFFLE Feng et al. (2023b), FedZeN Maritan et al. (2023), FedKSeed Qin et al. (2024),
FwdLLM Xu et al. (2024a), ZooPFL Lu et al. (2023b), FedMeZO Ling et al. (2024),wide leaf] ] ] [Adaptability
4.2), l1node, [Domain-Centric, l2node, [Domain-Adaptive Pre-Training, l3node, [  FMTDA Yao et al. (2022), FEDBFPT Wang et al. (2023) , leaf ] ] [Multi-Domain Adaptation, l3node, [  FedAPT Su et al. (2024), DiPrompT Bai et al. (2024b), leaf ] ] ] [Client-Centric, l2node, [Personalization, l3node, [  FedDAT Chen et al. (2024), Fed-MNMT Liu et al. (2023d),leaf ] ] [Client Clustering, l3node, [  FedLFC Guo et al. (2024b), FL-TAC Ping et al. (2024),leaf ] ] ] [System-Centric, l2node, [Resource-Heterogeneous, l3node, [  FedRA Su et al. (2023), HetLoRA Cho et al. (2024),leaf ] ] [ Split Learning, l3node, [ FedBERT Tian et al. (2022), FedSplitX Shin et al. (2023b),leaf] ] ] ] [Trustworthiness
4.3), l1node, [IP Protection, l2node, [Watermarking, l3node, [  WAFFLE Tekgul et al. (2021), DUW Yu et al. (2023b),leaf ] ] [Black-Box Tuning, l3node, [  Fed-BBPT Lin et al. (2023), pFedGPT Rui et al. (2024),leaf ] ] ] [Privacy Protection, l2node, [Privacy-Preserving Techniques, l3node, [  DP-FTRL Xu et al. (2023), DP-LoRA Liu et al. (2023c) , leaf ] ] [Privacy Attack, l3node, [  FILM Gupta et al. (2022), DRA Zhang et al. (2024c) , leaf ] ] ] [Attack Robustness, l2node, [Poisoning Attacks, l3node, [ Fed-EBD Li et al. (2024c), leaf ] ] [Defense Techniques, l3node, [ ClippedClustering Li et al. (2023b), Fed-FA Zhang et al. (2023d), leaf ] ] ] ] ]
Figure 1: Taxonomy of research in foundation models with federated learning.
4 Techniques
Recent work has begun to address challenges associated with adapting pre-trained FMs to specific downstream tasks in FL settings. In this section, we survey FM-FL techniques on three aspects, namely efficiency (Section 4.1), adaptability (Section 4.2), and trustworthiness (Section 4.3). As illustrated in Figure 1, we further refine them according to the key features of different methods.
4.1 Efficiency
There has been a considerable focus on developing resource-efficient approaches. This part describes techniques that improve resource efficiency.
4.1.1 Parameter-Efficient Fine-Tuning
Federated Parameter-Efficient Fine-Tuning (FedPEFT), originating from the fine-tuning practices of FMs Lester et al. (2021); Hu et al. (2022); Li and Liang (2021), is a suite of techniques designed to reduce both the computational load and the associated communication overheads Malaviya et al. (2023); Woisetschläger et al. (2024). In alignment with existing FM fine-tuning taxonomies Lialin et al. (2023); Ding et al. (2023), we present FedPEFT methods in three categories: selective methods, additive methods, and reparameterization-based methods.
Selective Methods
Selective methods fine-tune a small subset of the parameters, leaving the majority unchanged. In the field of LLMs, a prominent example of such methods is BitFit Ben Zaken et al. (2022), which only fine-tunes the bias terms. BitFit has inspired a series of studies in FedPEFT Bu et al. (2022); Sun et al. (2022a); Zhang et al. (2023f), demonstrating the superior communication efficiency of only updating the bias terms while still achieving competitive performance. More sophisticated methods strive to find sparse subnetworks for partial fine-tuning. Among them, various methods Seo et al. (2021); Li et al. (2021a); Tamirisa et al. (2024) advocate for the Lottery Ticket Hypothesis (LTH) Frankle and Carbin (2019), positing that a dense network contains many subnetworks whose inference capabilities are as accurate as that of the original network. FedSelect Tamirisa et al. (2024) is a representative method that encourages clients to find optimal subnetworks based on LTH and continually fine-tunes these derived subnetworks to encapsulate local knowledge. As another important aspect, RaFFM Yu et al. (2023c) proposes to prioritize specialized salient parameters by ranking them using salience evaluation metrics such as the $\ell_{1}$ and $\ell_{2}$ norms.
Additive Methods
Instead of fine-tuning a subset of model parameters, additive methods incorporate lightweight trainable blocks into frozen FMs and tune the additional parameters for model adaptation. These methods not only enhance computational and communicational efficiency but also introduce an extra benefit: personalization Lu et al. (2023a), i.e., the integration of these supplementary parameters allows for the customization of heterogeneous models tailored to specific local data characteristics or user preferences. Key branches within additive methods include adapter tuning and prompt tuning. Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models Houlsby et al. (2019); Hu et al. (2022). On the other hand, prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer Liu et al. (2023a); Dong et al. (2023). More details on these methods are provided in Appendix A.
Reparameterization-based Methods
The hypothesis behind reparameterization-based methods is that fine-tuning adaptations can be re-parameterized into optimization within low-rank subspaces Aghajanyan et al. (2021). Low-Rank Adaptation (LoRA) Hu et al. (2022), as a popular PEFT method from the area of LLMs, reduces the number of trainable parameters for downstream tasks by representing the weight updates with two smaller matrices (called update matrices) through low-rank decomposition Ding et al. (2023). When optimizing a parameter matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the update equation can be written as: $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$. The core idea of LoRA is to freeze the original matrix $\mathbf{W}$ while approximating the parameter update $\Delta\mathbf{W}$ by low-rank decomposition matrices, i.e., $\Delta\mathbf{W}=\mathbf{A}\cdot\mathbf{B}^{\top}$, where $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are the trainable parameters for task adaptation and $k\ll\min(m,n)$ is the reduced rank. The trainable parameter size is then reduced from $mn$ to $k(m+n)$. The major benefit of LoRA is that it can largely save memory and storage usage. A straightforward way to perform federated finetuning with LoRA is to train the LoRA modules $\mathbf{A}$ and $\mathbf{B}$ with homogeneous rank $k$ across all clients with standard FL such as FedAvg McMahan et al. (2017). Serval studies have shown that this method can achieve an outstanding level of trade-off between performance and communication overhead for a wide range of FMs, including language models Zhang et al. (2024b, 2023f), vision-language models Nguyen et al. (2024), and speech-to-text models Du et al. (2024).
Figure 2: Taxonomy of Federated Parameter-Efficient Fine-Tuning (FedPEFT). Apart from efficiency, some methods also account for other considerations, such as data and resource heterogeneity challenges that are identified in Section 3.2 and black-box tuning (see Section 4.3).
Comparison of FedPEFT methods
Figure 2 depicts the taxonomy of FedPEFT with representative methods. Note that some methods may belong to multiple overlapping categories. To compare the communication efficiency of different FedPEFT methods, Table 1 gives a brief overview of experimental evaluations from representative studies. Compared to full-model fine-tuning, FedPEFT methods only require 0.1%-30% communication overhead. We note that the differences can be attributed to several factors, including model complexity and implementation details.
4.1.2 Model Compression
Model compression refers to the techniques used to reduce the size of models, thereby improving resource efficiency Shah and Lau (2023).
Table 1: Comparison of Federated Parameter-Efficient Fine-Tuning (FedPEFT) Methods.
Category Representative Work Modality Model # Full Params. # Train. Params. Training Accel. Comm. Cost Selective RaFFM Yu et al. (2023c) Txt. BERT-Large (2019) 336M 100M 6.13×6.13$\times$6.13 × 29.8% FedBF Zhang et al. (2023f) Txt. Roberta-Base (2019) 125M 0.66M $1.6\%$ Additive Adapter FedAP Zhang et al. (2023f) Txt. Roberta-Base (2019) 125M 2M $1.6\%$ FedCLIP Lu et al. (2023a) Vis.-Txt. ViT-B/32 (2020a) 150M 0.53M 3.5% FedDAT Chen et al. (2024) Vis.-Txt. ALBEF (2021b) 290M 2.86M 9.9% C2A Kim et al. (2023) Txt. DistilBERT (2020) 66M 0.06M 0.1% Fed-MNMT Liu et al. (2023d) Txt. mBART-50 (2020) 611M 8M 1.3% AdaFL Cai et al. (2023) Txt. BERT (2019) 110M 0.61M 1.63×1.63$\times$1.63 × 0.6% Prompt PromptFL Guo et al. (2023) Vis.-Txt. ViT-B/16 (2021) 87M 0.87M 2.38×2.38$\times$2.38 × 0.9% MFPT Zhao et al. (2024b) Txt. XLM-RoBERTa (2020) 270M 1.2M 0.4% FedAPT Su et al. (2024) Vis.-Txt. ViT-B/32 (2020a) 88M 2.8M 3.2% FedSP Dong et al. (2023) Txt. GPT2-XL (2019) 1.6B  [^1] M 0.5% Reparameterization-based Methods SLoRA Babakniya et al. (2023a) Txt. DistilBERT (2020) 67M 0.7M 13.47×13.47$\times$13.47 × 5.8% LP-FL Jiang et al. (2023a) Txt. BERT-Large (2019) 336M 100M 30% FedMS Wu et al. (2023c) Vis.-Txt. ViT-B/16 (2021) 87M 8.6M 10% pFedS2T Du et al. (2024) Aud. Whisper (2023) 254M 10.1M 4% FFA-LoRA Sun et al. (2024b) Txt. RoBERTa-Large (2019) 355M 0.39M 0.1%
Sparsification
Model sparsification methods reduce communication burden by only transmitting a subset of FM parameters across the network Jiang et al. (2023c). Typical methods focus on identifying and cultivating high-potential subnetworks Frankle and Carbin (2019); Tsouvalas et al. (2023).
Quantization
Quantization is well-established in both the FM and FL domains Xu et al. (2024b); Reisizadeh et al. (2020), which involves decreasing the precision of floating-point parameters for mitigating the storage, computational, and communication demands. Quantization is orthogonal to other resource-efficient techniques, making it feasible to combine them for greater efficiency and flexibility Lit et al. (2022).
4.1.3 Zeroth-Order Optimization
In contrast to the use of gradient descent in most FL optimization algorithms, a particular line of research advocates for the removal of BackPropagation (BP) Malladi et al. (2023a) in favor of Zeroth-Order Optimization (ZOO) Fang et al. (2022); Li and Chen (2021). BP-free methods conserve memory needed for computing gradients and minimize communication overhead for model aggregation Qin et al. (2024), making FMs more accessible for lower-end devices, thereby enhancing their applicability in diverse hardware environments.
ZOO methods primarily rely on perturbation methods to estimate gradients with forward propagation. Given a model with parameters $\bm{\theta}\in\mathbb{R}^{d}$ and a loss function $\mathcal{L}$, a typical gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
^$\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta};\mathcal{B})}{2\epsilon}\bm{z},$ (2)
where $\bm{z}\in\mathbb{R}^{d}$ with $\bm{z}\sim\mathcal{N}(0,\bm{I}_{d})$ and $\epsilon$ is the perturbation scale  Duchi et al. (2015). It requires only two forward passes through the model to compute the estimation of gradient, serving as a memory-efficient alternative to BP. However, Eq. (2) provides a biased gradient estimation, leading to a certain degree of information loss Liu et al. (2020). Alternatively, many studies opt for two-point gradient estimators that can yield a more stable and reliable approximation Spall (1992); Malladi et al. (2023a); Lin et al. (2023); Ling et al. (2024). The standard two-point gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
^$\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon\bm{z};\mathcal{B})}{2\epsilon}\bm{z}.$ (3)
Based on the above gradient estimation frameworks, recent work, such as that by Xu et al. (2024a); Lu et al. (2023b), has initiated preliminary explorations into the deployment of both FedPEFT and full-model fine-tuning of billion-sized FMs, like LLaMA, on mobile devices. The naive ZOO methods remain impractical for training large FMs in standard FL frameworks such as FedAvg, as they still result in a significant communication burden for model aggregation. In light of this, FedKSeed Qin et al. (2024) was proposed to further reduce communication overheads between the server and clients by using just a few random seeds and scalar gradients, requiring only a few thousand bytes for communication.
Although ZOO methods have shown promise in resource-efficient FL Ling et al. (2024), they generally require many iterations to achieve strong performance Malladi et al. (2023b). Compared to the well-established BP-based optimization, ZOO is still in the early stages of development, particularly for FM-FL settings, necessitating further research and optimization.
4.2 Adaptability
Adaptation refers to the process of tailoring a pre-trained FM to perform effectively across varying FL settings and scenarios. This mainly includes the capability to learn from different domains, cater to individual user needs, and work across diverse devices while retaining overall performance and efficiency. We focus on three key aspects of adaptation, namely domain-centric adaptation, client-centric adaptation, and system-centric adaptation.
4.2.1 Domain-Centric Adaptation
Domain-centric adaptation focuses on adapting FMs within specific domains by addressing the domain diversity across client datasets.
Domain-Adaptive Pre-Training
Despite being heavily reliant on large-scale and public datasets for their initial training, FMs often require further Domain-Adaptive Pre-Training (DAPT) with domain-specific data for tasks that necessitate specialized knowledge Gururangan et al. (2020); Guo and Yu (2022). In domains like healthcare, FL allows for the continued pre-training of these models using sensitive, domain-specific data without compromising privacy. Based on this idea, Jiang et al. (2023b) proposed FFDAPT, a computational-efficient further pre-training algorithm that freezes a portion of consecutive layers while optimizing the rest of the layers. Similarly, Wang et al. (2023) proposed FEDBFPT that builds a local model for each client, progressively training the shallower layers of local models while sampling deeper layers, and aggregating trained parameters on a server to create the final global model.
Multi-Domain Adaptation
Given that client data may belong to various domains in real-world FL scenarios, some efforts Feng et al. (2023c); Su et al. (2024) have been devoted to facilitating multi-domain collaborative adaptation. Feng et al. (2023c) applied a pre-trained CLIP to the multi-domain scenario and proposed an adaptive prompt tuning method that uses domain-specific keys to generate prompts for each test sample. Furthermore, Su et al. (2024) employed knowledge distillation to selectively distill global knowledge based on an entropy measure, improving the generalization across different domains.
4.2.2 Client-Centric Adaptation
Client-centric adaptation refers to the process of tailoring an FM to meet the specific needs or preferences of individual clients while leveraging the decentralized and privacy-preserving nature of FL. Particularly, we discuss two types of popular personalized methods as follows:
Personalization
Adapter-based methods introduce small, trainable adapters into the frozen pre-trained FMs, allowing for client-specific model adaptation without altering the original FL. FedDAT Chen et al. (2024) leverages a dual-adapter structure, with personalized adapters focusing on client-specific knowledge and a global adapter maintaining client-agnostic knowledge. FedDAT executes bi-directional knowledge distillation between personalized adapters and the global adapter to regularize the client’s updates and prevent overfitting. Prompt-based methods involve using client-specific soft prompts to guide the model’s response. pFedPG Yang et al. (2023a) trains a prompt generator to exploit underlying client-specific characteristics and produce personalized prompts for each client, thereby enabling efficient and personalized adaptation.
Client Clustering
This branch of study aims to cluster clients based on the underlying relationships and tailor FMs for the client group with similar data distributions, thus reducing the negative impact of data heterogeneity and improving accuracy. Guo et al. (2024b) proposed a FedPEFT-based framework for multilingual modeling, which employs language family clustering to alleviate parameter conflicts of LoRA tuning.
4.2.3 System-Centric Adaptation
System-centric aims to improve adaptability at the system level. This involves handling resource heterogeneity in the FL systems while ensuring training efficiency and model utility.
Resource-Heterogeneous Methods
Cross-device FL systems may be composed of devices equipped with heterogeneous resources, leading to disparities where certain devices exhibit more efficient model training than others Chen et al. (2024). To address this issue, several methods have been developed to customize model architectures for resource-heterogeneous FL systems. In FL environments possessing heterogeneous resources, LoRA-based FedPEFT exhibits distinctive flexibility and adaptation in fine-tuning frozen FMs without overburdening client devices. Su et al. (2023) suggested assigning LoRA adapters to varying numbers of layers for heterogeneous clients according to a randomly generated mask matrix. An alternative and more targeted idea is to choose diverse LoRA ranks across clients based on their system capabilities. Bai et al. (2024a) proposed FlexLoRA to adjust local LoRA ranks dynamically. FlexLoRA reconstructs the uniform full-sized LoRA module $\Delta\mathbf{W}$ for server-side model aggregation followed by an SVD-based parameter redistribution. However, concurrent research by Cho et al. (2024) has empirically demonstrated that the reconstruct-redistribute method suffers from performance loss compared to homogeneous LoRA. Instead, they proposed HetLoRA Cho et al. (2024) that utilizes zero-padding to align module size before aggregation. It then truncates the global LoRA modules for the specific rank of the next selected clients.
Split Learning
Split learning addresses the resource heterogeneity between servers and clients by splitting a large model at a cut layer into client and server models Thapa et al. (2022). For each training step, the output tensor, so-called smashed data, from the client model and the corresponding labels are transmitted over to the server. The server continues the forward propagation by processing the smashed data through its remaining layers; it then computes the loss using the transmitted label and performs backpropagation. The gradient generated at the first layer of the server model is then transmitted back to the client for further backpropagation. Along this line, FedBERT Tian et al. (2022) proposes to leverage split learning for training the BERT model, showing the feasibility of training large FMs in FL settings. FedSplitX Shin et al. (2023b) is a more fine-grained method that allows multiple partition points for model splitting, accommodating more diverse client capabilities. Compared to conventional FL, split learning scales better with the size of FMs as it communicates only small-sized smashed data instead of model parameters Singh et al. (2019). Despite its merits, split learning is highly dependent on the network connection quality. Given that server-client interactions occur at every step of the optimization process Zheng et al. (2023), communication delays cause a more significant impact on efficiency.
4.3 Trustworthiness
This line of work aims to enhance trustworthiness throughout the FM-FL lifecycle, covering a variety of key aspects including, but not limited to, IP protection, privacy protection, and attack robustness.
4.3.1 IP Protection
Existing IP protection involves safeguarding ownership of FMs from unauthorized use (e.g., model theft) Tekgul et al. (2021). We discuss the following two mainstream IP protection strategies: watermarking and black-box tuning.
Watermarking
Watermarking is a well-known deterrence technology for model IP protection by providing the identities of model owners to demonstrate ownership of their models Adi et al. (2018). Tekgul et al. (2021) proposed WAFFLE, the first solution that addresses the ownership problem by injecting a watermark into the global model in FL environments. Recently, Yu et al. (2023b) proposed DUW that embeds a client-unique key into each client’s local model, aiming to identify the infringer of a leaked model while verifying the FL model’s ownership.
Black-Box Tuning
Black-Box Tuning (BBT) is a set of ZOO-based methods that fine-tune FMs without direct access to model parameters Sun et al. (2022c, b). BBT methods are often additive, introducing additional parameters while keeping the original model frozen (see Section 4.1.1). Fed-BBPT Lin et al. (2023) is a general prompt tuning framework that facilitates the joint training of a global lightweight prompt generator across multiple clients. FedBPT Sun et al. (2024a) adopts a classic evolutionary-based ZOO method, CMA-ES Hansen and Ostermeier (2001), for training an optimal prompt that improves the performance of frozen FMs. ZooPFL Lu et al. (2023b), on the other hand, applies coordinate-wise gradient estimate to learn input surgery that incorporates client-specific embeddings. BBT allows for local fine-tuning of FMs while not infringing IP constraints. However, current research in this line is limited to few-shot learning with small datasets for LLM fine-tuning Sun et al. (2022b), while larger datasets and other modalities remain unexplored.
4.3.2 Privacy Protection
Protecting privacy in FM-FL requires both designing protective measures and studying privacy attack strategies.
Privacy-Preserving Techniques
Differential Privacy (DP) is a theoretical framework that governs privacy boundaries and manages the tradeoff between privacy and model convergence Wei et al. (2020); Xu et al. (2023). DP-based FL approaches often add artificial noise (e.g., Gaussian noise) to parameters at the clients’ side before aggregating to prevent information leakage Xu et al. (2023). Besides, DP is compatible with most FedPEFT methods. For instance, Sun et al. (2024b) showed that DP noise can even be amplified by the locally “semi-quadratic” nature of LoRA-based methods, motivating the integration of LoRA with DP to improve resource efficiency while maintaining data privacy Liu et al. (2023c). In addition to DP, Secure Multi-Party Computation (SMPC) Mugunthan et al. (2019) and Homomorphic Encryption (HE) Zhang et al. (2020) are also effective privacy-preserving mechanisms. However, they do not scale well enough for large-scale deployments in FM-FL.
Privacy Attack
Privacy attacks in FM-FL involve extracting sensitive information from the data used in training, even though the data itself is not directly shared. Major attacks include membership inference attack and data reconstruction attack, where the former aims to determine whether a specific data sample is in a victim client’s training set, and the latter strives to reconstruct original input data from the model parameters or gradients Ren et al. (2024). Regarding membership inference attacks, Vu et al. (2024) revealed the vulnerabilities of popular LLMs, including BERT, DistilBERT, and OpenAI’s GPTs. In terms of data reconstruction attacks, Gupta et al. (2022) presented an attack FILM, which recovers private text data by extracting information from gradients transmitted during training despite employing a DP mechanism.
4.3.3 Attack Robustness
Due to the distributed characteristic of optimization, FL is vulnerable to poisoning attacks Lyu et al. (2022); Rodríguez-Barroso et al. (2023), wherein certain participants may deviate from the prescribed update protocol and upload arbitrary parameters to the central server.
Poisoning Attacks
Depending on the adversarial goals, poisoning attacks in FL can be classified as targeted and untargeted Jere et al. (2020). Targeted attacks, like backdoor attacks, aim to manipulate the global model to generate attacker-desired misclassifications for some particular samples Xie et al. (2020); Bagdasaryan et al. (2020). In contrast, untargeted attacks seek to degrade the model’s overall performance indiscriminately Fang et al. (2020). In addition to the well-recognized attacks on conventional FL studies Li et al. (2023b, 2024b), FM-FL also faces potential threats from compromised pre-trained FMs Li et al. (2023c). Thus, The attacker can introduce backdoors to downstream tasks without prior knowledge Shen et al. (2021). Specifically, Li et al. (2023d) proposed Fed-EBD that introduces a backdoor-compromised FM to generate a public, synthetic dataset for FL training. The clients’ models, pre-trained on this dataset, inherit the backdoor throughout the training.
Defense Techniques
As for defenses, robust aggregation rules are widely applied to make an attack-resilient estimation of the true updates and exclude the influence of malicious updates Blanchard et al. (2017); Yin et al. (2018); Chen et al. (2017); Li et al. (2023a). Other research directions include trust-based strategies Cao et al. (2021); Xu et al. (2022); Park et al. (2021) and variance-reduced algorithms Gorbunov et al. (2023); Wu et al. (2020b). Although these techniques have been widely examined in various FL settings, their effectiveness has yet to be explored in the FM-FL paradigm.
Table 2: A list of representative studies on the applications of FM-FL. Abbreviations: LoRA Tuning (LT), Adapter Tuning (AT), Full-Parameter Tuning (FT), Selective Tuning (ST), Prompt Tuning (PT).
Domain/Application Task Representative Work On-Device Personalization Modality Backbone Fine-Tuning Multilingual NLP Language Understanding FedKC Wang et al. (2022) Txt. mBERT FT Multi-Tasks PMMFL Weller et al. (2022) Txt. mBERT FT Machine Translation Fed-MNMT Liu et al. (2023d) Txt. mBART-50 AT Machine Translation FL-MetaSend Chu et al. (2024) Txt. M2M-100 ST Multi-Tasks MFPT Zhao et al. (2024b) Txt. XLM-RoBERTa PT Speech Speech-to-Text pFedS2T Du et al. (2024) Aud. Conformer/Whisper LT Speech Recognition FedASR Jia et al. (2023) Aud. RNN-T AT Speech Recognition FedE2EASRAzam et al. (2023a) Aud. CTC-AED FT Recommendation General PPLR Zhao et al. (2024a) Txt. LLaMA-7B/LongFormer FT General TransFR Zhang et al. (2024a) Txt. DistBERT AT General GPT-FedRec Zeng et al. (2024) Txt. ChatGPT NA Healthcare Mental Health Prediction FedTherapist Shin et al. (2023a) Txt. BERT & LLaMa-7B LT MRI Reconstruction FedPR Feng et al. (2023a) Vis. Swin Transformers PT
5 Applications of FM-FL
In this part, we briefly review the recent progress on FM-FL applications. Table 2 lists representative work on specific applications and domains.
5.1 FM-FL for Multilingual NLP
Multilingual NLP refers to the techniques that handle multiple natural languages Pires et al. (2019), often to perform equally well across them Wu and Dredze (2020). Earlier research Johnson et al. (2017) has shown that parameter sharing among different languages boosts the model’s performance in multilingual NLP, especially for low-resource languages for which significantly less content is available. However, real-world multilingual text data is often distributed across devices or regions, with each client (user) accessing only a limited subset of languages, where transferring the data to a central server is often problematic or prohibited due to privacy issues Wang et al. (2022). Thanks to its inherent privacy-preserving characteristic, FL holds promise in breaking the barriers of cross-lingual modeling and data isolation by allowing models to learn from decentralized datasets.
The pioneer work by Weller et al. (2022) has firstly demonstrated that fine-tuning pre-trained language models with FL can perform similarly to pre-trained models fine-tuned with the standard centralized method under multilingual NLP settings. Various subsequent studies have focused on adapting pre-trained FMs through FedPEFT techniques such as adapter tuning Liu et al. (2023d), prompt tuning Zhao et al. (2024b), and LoRA Guo et al. (2024b), aiming to enhance training efficiency.
Considering the adverse effect of conflicting parameters from diverse languages during federated fine-tuning, recent studies have exploited clustering strategies to alleviate this issue. For instance, Wang et al. (2022) applied $k$-means clustering on each client’s data to obtain representative knowledge, specifically the clustered data centroids. These centroids were then shared across clients for local training, enriching training data and addressing the challenges associated with data heterogeneity. Another compelling strategy along this line is language family-based clustering. Liu et al. (2023d) explored various clustering strategies to group adapter parameters to mitigate the negative effects of multilingual data heterogeneity, showing that language family-based clustering significantly outperforms the other clustering strategies. Similarly, Guo et al. (2024b) proposed fine-tuning FMs with LoRA and language family-based clustering to address the heterogeneity issue of multilingual modeling.
General downstream tasks include language modeling Wang et al. (2022), machine translation Liu et al. (2023d); Chu et al. (2024), and text classification Weller et al. (2022). In addition, some studies also focus on more specific applications such as medical transcript analysis Manoel et al. (2023) and hate speech detection Akshay and Rahul (2024). These advancements illustrate the applicability of FM-FL across a wide range of scenarios in multilingual NLP.
5.2 FM-FL for Speech
With the development of AI, researchers have also carried out many studies on speech-related FMs, e.g., wav2vec 2.0 Baevski et al. (2020) and Whisper Radford et al. (2023). In this field, the adaptation of FMs often relies on FL to facilitate scenarios where the audio data is privacy-sensitive. Compared to other data modalities, speech-related FM-FL applications especially attract excessive attention to the aspects of on-device training and personalization, motivated by the following considerations: (1) Audio data is continually generated on end-devices such as mobile phones, and owned by individual users—thus it should be processed locally, rather than being transferred elsewhere; (2) Although FL takes advantage of all user data to collectively train one model that maximizes speaker-independent accuracy, such a one-model-fits-all solution can be sub-optimal for individual users Jia et al. (2023). Specific tasks in this field include Automatic Speech Recognition (ASR) Azam et al. (2023b) and Speech-to-Text (S2T) Du et al. (2024).
5.3 FM-FL for Recommendation
Federated Recommendation (FR) strives to capture underlying user preferences and recommend appropriate information to users while safeguarding data privacy Bobadilla et al. (2013); Zhang et al. (2023a). Typical FR systems consist of a server and multiple clients, where clients represent individual users or local data servers possessing smaller datasets and retaining private user information Ammad-Ud-Din et al. (2019). These clients collaborate to train a global model while ensuring their data privacy protection by abstaining from direct data sharing Zeng et al. (2024); Zhang et al. (2023a). Recently, LLM-based recommendations have been gaining increasing attention Wu et al. (2023b) due to their strong capacities in language understanding and domain generalization. The benefits are mainly twofold: (1) LLMs mitigate the cold-start issue by utilizing textual descriptions to make recommendations without the need for extensive historical data Zhang et al. (2023c); (2) The inherent transferability of LLMs allows them to apply cross-domain knowledge and side information to improve accuracy and relevance across diverse items and user interests Gao et al. (2023).
One straightforward way to adapt FMs for FR is by fine-tuning them with historical user-item data. More specifically, FedPEFT techniques such as adapter tuning Zhang et al. (2024a) and split learning Zhao et al. (2024a) can be employed to improve resource efficiency. Apart from parameter fine-tuning, LLMs can also be adapted to assist the recommendation in a zero-shot paradigm through prompt engineering (i.e., without parameter tuning) Gao et al. (2023). For example, Zeng et al. (2024) proposed GPT-FedRec, a two-stage FR framework that leverages ChatGPT for its powerful zero-shot generalization ability. Firstly, GPT-FedRec facilitates hybrid retrieval by collaboratively training ID and text retrievers, after which the retrieved results are transformed into text prompts and submitted to GPT for re-ranking in the second stage. Additionally, Guo et al. (2024a) employed a pre-trained BERT to obtain the representation vectors of item descriptions, which are then fed into a recommender system as augmented input.
5.4 FM-FL for Healthcare
FMs, especially LLMs, have been found to excel in healthcare applications, showcasing impressive capabilities in tasks like mental health analysis Yang et al. (2023b), disease diagnosis Panagoulias et al. (2024), and drug discovery Chenthamarakshan et al. (2023). However, it raises privacy concerns to upload the health information of patients Tang et al. (2023) into a commercial server that supports the FMs. Meanwhile, FL has consistently received widespread attention in the healthcare domain Lincy and Kowshalya (2020); Rieke et al. (2020); Joshi et al. (2022), driven by the need for collaborative model training across different medical institutions without compromising patient data privacy. By breaking the barriers of private data availability, the FM-FL paradigm shows the potential to further harness the power of FMs in the healthcare domain.
A recent study Shin et al. (2023a) presents a mobile mental health monitoring system, FedTherapist, which leverages user speech and keyboard input to fine-tune FMs with FL, demonstrating superior accuracy in mental health prediction tasks such as depression, stress, and mood prediction. Another representative study Feng et al. (2023a) focuses on Magnetic Resonance Imaging (MRI) reconstruction, which involves retrieving a complex-valued image from its under-sampled signal. The authors adopted an FM pre-trained on public datasets and trained visual prompts from decentralized clinical datasets via a personalized FL mechanism, thereby reducing communication costs and achieving competitive performance on limited local data.
Despite the efforts, it has been shown that FMs in healthcare risk generating misleading information due to their imperfect understanding of complex medical data Jeblick et al. (2024).
6 Future Directions
Although recent work has already begun to address the challenges discussed in Section 3.2, many critical open directions are yet to be explored. Here, we outline several representative ones.
Multimodal FM-FL
With the development of mobile technology and IoT infrastructures Brunete et al. (2021), numerous edge devices produce data from a range of modalities, such as sensory, visual, and audio. In the era of FMs, the success of LLMs and their multimodal derivatives Ramesh et al. (2021); Google (2023); OpenAI (2024) have demonstrated the potential of multimodal FMs. The potential opportunities and challenges for multimodal FM-FL have yet to be explored.
Continual Learning
Continual learning enables models to adapt to new data over time, improving their performance and accuracy. By incorporating new data into the model training process, FL and FMs can continuously improve and adapt to changing environments and user needs Yang et al. (2024a). Future directions may involve leveraging transfer learning techniques in continual learning for FL and FMs. Models can transfer knowledge from previous tasks or domains to new ones, enabling more efficient adaptation Good et al. (2023).
Efficient Federated Black-Box Tuning
In scenarios where gradient access is unavailable, preliminary efforts have focused on federated fine-tuning black-box FMs Lin et al. (2023); Sun et al. (2024a); Lu et al. (2023b); Rui et al. (2024) utilizing ZOO. However, ZOO’s noticeably slower convergence rates, especially in high-dimensional contexts compared to gradient-based methods Golovin et al. (2020), indicate an important direction for further research. The impact of these slower convergence rates on overall efficiency and computational load within FL, particularly concerning large-scale FMs, has not been adequately investigated and understood.
FL with AI-Generated Content
AI-Generated Content (AIGC) denotes content produced via advanced generative FMs Wu et al. (2023a). The strong generative capability of FMs offers the advantage of rapidly automating the creation of inexhaustible synthetic data. This capability positions AIGC as a valuable supplementary data source for model training and evaluation in many tasks Xu et al. (2024c). Despite some efforts Zhang et al. (2023b), more potential opportunities and challenges for AIGC-aided FL have yet to be explored.
7 Conclusions
In this survey, we have meticulously surveyed the intersection of FM and FL. We identified core challenges in efficiency, adaptability, and trustworthiness and proposed a comprehensive taxonomy of techniques in response to these challenges. In addition, we discussed future directions and applications in this research field, hoping to attract more breakthroughs in future research.
Limitations
FM and FL are very fast-moving fields. We have put a lot of effort into including the latest research efforts in the community in this survey. Therefore, we believe that our survey will help to inspire and push further research and innovation in these important areas. Our survey does not focus on experimental evaluation of the available ideas and systems. We believe that would be an important next step that we are leaving for future work.
Appendix A Additional Details of Adapter Tuning
A.1 Adapter Tuning
Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models Houlsby et al. (2019); Hu et al. (2022). A straightforward implementation of adapter tuning is to collaboratively train a shared adapter among all clients in the FedAvg manner, as highlighted by Sun et al. (2022a). Based on FedAvg, FedCLIP Lu et al. (2023a) incorporates an attention-based adapter for the image encoder in CLIP models Radford et al. (2021). In the domain of multilingual machine translation, where different language pairs exhibit substantial discrepancies in data distributions, Fed-MNMT Liu et al. (2023d) explores clustering strategies that group adapter parameters and makes inner-cluster parameters aggregation for alleviating the undesirable effect of data discrepancy. Another representative approach named C2A Kim et al. (2023) employs hypernetworks Ha et al. (2017) to generate client-specific adapters by conditioning on the client’s information, maximizing the utility of shared model parameters while minimizing the divergence caused by data heterogeneity.
A.2 Prompt Tuning
Prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer Liu et al. (2023a); Dong et al. (2023). Compared to full fine-tuning, it achieves comparable performance but with 1000×1000$\times$1000 × less parameter storage and communication Jia et al. (2022). A variation of prompt tuning, FedPerfix Sun et al. (2023) uses a local adapter to generate the prefixes and aggregate the original self-attention layers.
Depending on target modalities, prompt tuning in current literature can be further classified into three categories:
  • Textual Prompt Tuning. Task-specific prompt embeddings are combined with the input text embeddings, which are subsequently fed into language models. These soft prompts serve as instructive contexts to influence the generation process of LLMs by steering the probability distribution of the next token Dong et al. (2023).
  • Visual Prompt Tuning. Taking inspiration from advances in efficiently tuning LLMs, prompts are also introduced in the input space of vision models Jia et al. (2022). Naive implementations introduce prompts at the pixel level, acting as a form of data augmentation Li et al. (2024a). Alternatively, one could also insert the prompts as latent vectors for the first Transformer layer Deng et al. (2024); Yang et al. (2023a). Nevertheless, an empirical study Jia et al. (2022) has suggested that it is easier for visual prompts to learn condensed task-dependent signals in the latent input space of Transformers.
  • Textual-Visual Prompt Tuning. Unlike single-modal FMs, vision-language FMs can process and interpret both visual data and textual information, endowing them with powerful representation ability and transferability Radford et al. (2021). Based on vision-language FMs like CLIP, textual-visual prompt tuning shows promising capabilities in FL Guo et al. (2023), especially in cross-domain scenarios, where the model needs to generalize across varied domains and unseen classes Qiu et al. (2024).
Table 3: A list of existing FM-FL libraries and benchmarks. Missing or inapplicable details denoted by N/A. ✓ denotes a strong focus or presence; ✗ indicates no focus or absence; ◐ signifies a moderate focus or partial inclusion.
Library/Benchmark FL Backend LLM Support MultiModal FM Support FedPEFT On-Device Training Distributed & Clustered Differential Privacy Description FederatedScope-LLM Kuang et al. (2023) FederatedScope An end-to-end benchmark for efficient fine-tuning LLMs with FL NVIDIA FLARE Roth et al. (2024) NVFlare Scalable and efficient fine-tuning LLMs with FL FATE-LLM Fan et al. (2023) FATE Focuses on IP and privacy protection in federated LLM FedLLM FedML (2023) FedML An MLOps-supported training pipeline based on FedML OpenFedLLM Ye et al. (2024) N/A N/A An LLM framework focusing on FL instruction tuning/alignment Shepherd Zhang et al. (2024b) N/A Federated instruction tuning based on Hugging Face FedPETuning Zhang et al. (2023f) FedLab A benchmark comprising four FedPEFT methods FedLegal Zhang et al. (2023e) FedLab A benchmark comprising six legal NLP tasks under FL settings
Appendix B Libraries and Benchmarks
This part briefly introduces a series of available libraries and benchmarks for developing and examining FM-FL techniques. An overview is provided in Table 3.
  • FederatedScope-LLM Kuang et al. (2023) is an open-source package for fine-tuning LLMs via FL. Built on top of a popular FL backend FederatedScope Xie et al. (2023), it supports federated fine-tuning of LLMs under various FL scenarios, including FedPEFT and model personalization.
  • NVIDIA FLARE Roth et al. (2024) is an FL framework that allows researchers and data scientists to seamlessly move their machine learning and deep learning workflows into a federated paradigm.
  • FATE-LLM Fan et al. (2023) is an industrial-grade FL framework for LLM. Apart from FedPEFT, it provides a privacy hub integrating several IP protection and privacy-preserving mechanisms to protect model security and data privacy.
  • FedLLM FedML (2023) is an MLOps-supported training pipeline built upon the FedML AI platform He et al. (2020). FedLLM is compatible with popular LLM libraries such as HuggingFace and DeepSpeed to support a large range of FMs and datasets.
  • OpenFedLLM Ye et al. (2024) is a federated tuning framework for LLMs, which covers applications of instruction tuning and value alignment, diverse FL baselines, training datasets, and evaluation datasets.
  • Shepherd Zhang et al. (2024b) is a lightweight federated tuning framework. The local training process of Shepherd is built upon the implementations of Alpaca-LoRA Wang (2023), and Hugging Face’s PEFT Mangrulkar et al. (2022), enabling efficient fine-tuning.
  • FedPETuning Zhang et al. (2023f) is a pioneering federated benchmark for four representative FedPEFT methods, covering adapter tuning, prefix tuning, LoRA, and BitFit.
  • FedLegal Zhang et al. (2023e) is the very first real-world FL benchmark for legal NLP, which comprises five legal NLP tasks and one privacy task based on the data from Chinese courts.