### v2

# Synergizing Foundation Models and Federated Learning: A Survey

Abstract.
Over the past few years, the landscape of Artificial Intelligence (AI) has been reshaped by the emergence of Foundation Models (FMs). Pre-trained on massive datasets, these models exhibit exceptional performance across diverse downstream tasks through adaptation techniques like fine-tuning and prompt learning. More recently, the synergy of FMs and Federated Learning (FL) has emerged as a promising paradigm, often termed Federated Foundation Models (FedFM), allowing for collaborative model adaptation while preserving data privacy. This survey paper provides a systematic review of the current state of the art in FedFM, offering insights and guidance into the evolving landscape. Specifically, we present a comprehensive multi-tiered taxonomy based on three major dimensions, namely efficiency, adaptability, and trustworthiness. To facilitate practical implementation and experimental research, we undertake a thorough review of existing libraries and benchmarks. Furthermore, we discuss the diverse real-world applications of this paradigm across multiple domains. Finally, we outline promising research directions to foster future advancements in FedFM. Overall, this survey serves as a resource for researchers and practitioners, offering a thorough understanding of FedFM’s role in revolutionizing privacy-preserving AI and pointing toward future innovations in this promising area.
copyright: noneccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paperccs: Do Not Use This Code Generate the Correct Terms for Your Paper
1. Introduction
A significant paradigm shift in Artificial Intelligence (AI) has been catalyzed by the rapid advancement of Foundation Models (FMs) (Bommasani et al., 2021), such as BERT (Devlin et al., 2019), GPT series (Brown and others, 2020; OpenAI, 2022, 2024), and LLaMA series (Touvron et al., 2023a, b) in Natural Language Processing (NLP); ViTs (Dosovitskiy et al., 2021) and SAM (Kirillov et al., 2023) in Computer Vision (CV); CLIP (Radford et al., 2021), DALL-E (Ramesh et al., 2021), Gemini (Google, 2023), and GPT-4o in multimodal applications. These FMs have become pivotal in a myriad of AI applications across diverse domains. The remarkable generalization capabilities of FMs are attributed to large-scale pre-training (Gunasekar et al., 2023), which enables them to learn robust representations of language, vision, and multimodal data. To leverage these representations for specific applications, adaptation techniques such as fine-tuning are widely adopted to tailor FMs to diverse downstream tasks (Zhao et al., 2023).
While general-purpose FMs can leverage openly accessible data from the Internet, domain-specific FMs require proprietary data. It is, however, challenging to collect vast amounts of proprietary data and perform centralized pre-training or fine-tuning for domain-specific models, due to privacy restrictions (Jo and Gebru, 2020), i.e., General Data Protection Regulation (GDPR) (GDPR, 2016) and California Consumer Privacy Act (CCPA) (CCPA, 2023). Particularly in domains such as law, healthcare, and finance, where data is inherently privacy-sensitive, there is a pressing need for stringent privacy safeguards. Furthermore, given that data often constitutes a pivotal asset for enterprises, its widespread distribution is prohibitive. Consequently, there is an urgent need for novel strategies to handle data availability and facilitate model training, thereby unlocking the potential of domain-specific applications of FMs whilst respecting data privacy.
To address the challenges of data privacy in model training, Federated Learning (FL) (McMahan et al., 2017) has emerged as a promising paradigm. FL facilitates collaborative model training across decentralized clients without sharing raw data, thereby preserving privacy. Specifically, FL involves periodic interactions between the server and decentralized clients to exchange trainable model parameters without requiring private client data. Recognizing such a benefit, integrating FMs with FL, often termed Federated Foundation Models (FedFM), presents a compelling solution for domain-specific FMs (Zhuang et al., 2023; Yu et al., 2024).
Table 1. Summary of previous surveys related to FedFM (: Included, : Partially included, : Not included)
Survey Year Taxonomy Technical Focus Library & Benchmark Application Future Direction
Efficiency Adaptability Trustworthiness
Zhuang et al. (2023) (2023)
Chen et al. (2023) (2023)
Yu et al. (2024) (2024)
Woisetschläger et al. (2024) (2024)
Hu et al. (2024) (2024)
Li and Wang (2024) (2024)
Yao et al. (2024b) (2024b)
Kang et al. (2025) (2025)
Wu et al. (2025b) (2025b)
Fan et al. (2025c) (2025c)
Ren et al. (2025) (2025)
Ours 2026
Driven by the surging interest in the intersection of FMs and FL, several surveys have investigated various dimensions of the FedFM paradigm. Table 1 compares our work with prior surveys in terms of coverage of key aspects. Most existing surveys focus on presenting the vision and challenges, but few analyze the overall system performance from multiple perspectives. Specifically, early contributions primarily articulated the motivation and outlined the key challenges of integrating FMs with FL (Zhuang et al., 2023; Yu et al., 2024; Chen et al., 2023). Subsequent work has considered specific topics but mostly focuses on efficiency (Woisetschläger et al., 2024; Hu et al., 2024; Fan et al., 2025c; Yao et al., 2024b), with very few studies on adaptability and trustworthiness (Kang et al., 2025; Wu et al., 2025b; Li and Wang, 2024). Moreover, these works rarely provide a unified taxonomy that captures the full landscape. Also, none of them considers all three important metrics on efficiency, adaptability, and trustworthiness holistically. To the best of our knowledge, the work by Ren et al. (2025) represents the only prior attempt to provide a comparatively comprehensive taxonomy across various dimensions of FedFM. However, this work does not consider adaptability a critical dimension; instead, it structures its taxonomy around the conventional FL lifecycle. We argue that this emerging paradigm necessitates a novel perspective to fit new scenarios and demands of FMs.
Recognizing gaps in the existing literature, we conduct a systematic review of FedFM research to provide a thorough exploration of system performance in the integration of FMs and FL. More specifically, we provide a comprehensive study and analysis of the existing work across three major performance perspectives: efficiency, adaptability, and trustworthiness. Our main contributions are detailed as follows:
  • We present a novel multi-tiered taxonomy for FedFM. This framework systematically classifies existing techniques based on their characteristics and modeling approaches. It serves as a foundation for FedFM methodologies, guiding the design of future solutions.
  • We discuss publicly available libraries and benchmarks, including their attributes, to make it easy for readers to get started with their research.
  • We introduce typical applications of FedFM in critical domains like NLP, Recommendation, and Healthcare.
The remainder of this survey is organized as follows: Section 2 introduces background on FMs and FL. Section 3 presents the motivation and challenges for synergizing FMs and FL. Section 4, 5, and 6 review and summarize the recent progress from the three aspects of efficiency, adaptation, and trustworthiness. Section 8 explores the applications across various domains. Before concluding, we discuss representative future directions in Section 9.
2. Background
This section briefly describes the background of FMs and FL, respectively.
2.1. Foundation Models
An FM is a model that can be adapted to a wide array of tasks through fine-tuning after initial pre-training (Bommasani et al., 2021). The lifecycle of FMs typically involves pre-training on extensive generic data to establish the basis of their abilities (Bubeck et al., 2023), followed by adaptation to downstream tasks such as domain-specific question answering (Zhang et al., 2023d), and ultimately application in various domains.
FMs have sparked a significant paradigm shift in various fields of AI such as NLP, CV, speech and acoustics, and beyond. In the realm of NLP, the most prominent example is Large Language Models (LLMs) with substantial parameter sizes (Zhao et al., 2023). These models, such as ChatGPT and GPT-4 (OpenAI, 2022, 2024), demonstrate exceptional abilities in natural language understanding and generation, enabling them to comprehend and respond to user inputs with remarkable contextual relevance. This capability proves invaluable in applications like customer service, virtual assistants, and chatbots, where effective communication is paramount. Moreover, LLMs eliminate the need for training models from scratch for specific tasks, be it machine translation, document summarization, text generation, or other language-related tasks.
[…]
2.2. Federated Learning
FL (McMahan et al., 2017) is a learning paradigm that enables a collection of clients to collaboratively learn a shared global model by leveraging their private datasets in a distributed manner, assisted by the coordination of a central server. The general goal of FL is to find a parameter set $\bm{\theta}$ that minimizes the following distributed optimization objective:
(1) $\min\limits_{\bm{\theta}}F(\bm{\theta}):=\frac{1}{K}\sum_{k\in[K]}F_{k}(\bm{\theta}),$
where $K$ represents the total number of clients and $F_{k}(\bm{\theta})=\mathbb{E}_{\bm{z}\sim\mathcal{D}_{k}}[\ell(\bm{\theta};\bm{z})]$ denotes the expected risk of the $k$-th client. Here, $\mathcal{D}_{k}$ is the data distribution for the $k$-th client, and $\ell(\cdot;\cdot)$ is a user-specified loss function.
The most representative algorithms in the FL literature are the FedAvg-family algorithms (McMahan et al., 2017; Reddi et al., 2021). The standard FedAvg involves periodic interactions between the server and decentralized clients to exchange trainable model parameters. In this process, each client independently trains the model on its local data and sends the model updates to a central server. The server aggregates these updates by computing their average to update the global model, which is subsequently redistributed to the clients for further iterations. Many variants have been proposed to tackle issues such as convergence and local data heterogeneity (Diao et al., 2021). For example, FedProx (Li et al., 2020) and FedDyn (Acar et al., 2021) introduce regularizer terms to penalize client updates that are far away from the server model. A general framework FedOpt (Reddi et al., 2021) unifies adaptive optimizers (Adam, Yogi, etc.) and demonstrates superior convergence speed when compared to the naive FedAvg.
FL offers an efficient privacy-preserving way to train models on large-scale and diverse data (Kairouz and others, 2021), leading to its application across various domains such as healthcare (Lincy and Kowshalya, 2020; Rieke et al., 2020; Joshi et al., 2022), finance (Chatterjee et al., 2023; Liu et al., 2023b), and smart cities (Ramu et al., 2022; Pandya et al., 2023).
3. FedFM: Motivation & Challenges
In this section, we first motivate the synergy of FMs and FL (Section 3.1), then summarize the key challenges (Section 3.2).
3.1. Motivation
The integration of FMs and FL represents a compelling collaboration that leverages each other’s strengths to address their respective limitations, embodying a complementary relationship (Zhuang et al., 2023; Li and Wang, 2024).
FL expands data availability for FMs
By leveraging data from a wide range of sources in a privacy-preserving manner, FL makes it possible to build models on sensitive data in specific domains, such as healthcare (Lincy and Kowshalya, 2020; Joshi et al., 2022; Rieke et al., 2020) and finance (Chatterjee et al., 2023; Liu et al., 2023b). This enhances the diversity and volume of training data, improving model robustness and adaptability. Moreover, FL enables the integration of personal and task-specific data, allowing FMs to be customized for personal applications. For instance, Google has trained next-word-prediction language models on mobile keyboard input data with FL to improve user experience (Xu et al., 2023; Bonawitz et al., 2021).
FMs boost FL with feature representation and few-shot learning capabilities
By pre-training on large-scale generic data, FMs acquire essential knowledge and understanding capabilities (Brown and others, 2020), providing multiple benefits to FL. Firstly, they benefit FL systems by offering advanced feature representations and learning capabilities from the outset. Secondly, leveraging the pre-learned knowledge of FMs can accelerate the FL process, enabling efficient and effective adaptation to specific tasks with minimal additional training. Thirdly, FMs’ powerful generative capabilities could help FL overcome the data heterogeneity challenge by synthesizing extra data, thus accelerating model convergence (Huang et al., 2024b).
3.2. Core Challenges
In this section, we discuss challenges arising from FedFM across three major performance perspectives: efficiency, adaptability, and trustworthiness. They are considered important metrics for training and running the FMs across diverse downstream applications. We will explain them one by one below.
Efficiency Challenges
Efficiency challenges stem from the mismatch between the significant resource demands of FM training and the limited, heterogeneous system resources (e.g., mobile devices) within FL systems, such as communication bandwidth, computational power, and memory (Su et al., 2024a). The communication bottleneck of FL is induced by frequently exchanging training information between the server and clients over limited bandwidth channels (Kairouz and others, 2021). The substantial number of parameters in FMs further exacerbates this burden, thus hindering the training process.
Adaptability Challenges
Adaptability challenges arise from the adaptation of an FM to a specific downstream task (e.g., by fine-tuning) in FL settings. Key challenges include data heterogeneity and resource heterogeneity. Performance degradation in FL, attributed to heterogeneous data distributions among clients, is a well-recognized issue (Kairouz and others, 2021; Li et al., 2022). A recent study (Babakniya et al., 2023) has shown that such a performance penalty is even more substantial when fine-tuning FMs. For NLP tasks, data heterogeneity can manifest as variations in language, style, topic, or sentiment across datasets held by different clients. In multi-modal scenarios, the challenge is even more pronounced due to the inherent diversity in data types (e.g., text, images, and audio) (Yu et al., 2023a). Addressing data heterogeneity involves not just identifying and measuring it but also developing algorithms that are robust to such diversity, ensuring that the model can learn effectively from varied data contributions without compromising on performance. In terms of resource heterogeneity, the memory and computational resources of the devices for different participants may be diverse (Diao et al., 2021), which could cause delays for model synchronization and inactivation of some participants, i.e., stragglers, making it challenging to leverage the full potential of FMs in FL settings.
Trustworthiness Challenges
Trustworthiness challenges emphasize the concerns regarding privacy, security, and ethical considerations in the lifecycle of FedFM, from the pre-training and model adaptation to the application stages. We present two representative challenges from this perspective: (1) intellectual property: Intellectual Property (IP) protection in FedFM primarily involves attributing ownership rights for both models and data. From the server’s perspective, broadcasting a pre-trained model to multiple nodes for fine-tuning poses IP protection and security risks (e.g., model theft), necessitating measures to safeguard IP rights and ensure model integrity (Kang et al., 2025); (2) privacy leakage: Although FL does not immediately share data, studies have shown that it may not always guarantee sufficient privacy preservation (Geiping et al., 2020), as model parameters (e.g., weights or gradients) may leak sensitive information to malicious adversaries (Zhu et al., 2019). (3)Poisoning Attacks: FL systems are inherently vulnerable to attacks due to their wide attack surface and reliance on network communication (Li et al., 2024c). Poisoning attacks are carried out by malicious participants, aiming to bias the global model to the desire of attackers.
In the following sections (Sections 4-6), we will review and discuss the existing work with respect to efficiency, adaptability, and trustworthiness.
{forest}
Figure 1. Taxonomy of research in foundation models with federated learning.
4. Efficiency
The remarkable capabilities of FMs are often accompanied by substantial resource demands, posing significant challenges regarding efficiency of FedFM. To address these bottlenecks, research has increasingly prioritized resource-efficient approaches. In this section, we review and discuss the current state of the art in efficiency, focusing on three primary streams: parameter-efficient fine-tuning, model compression, and zeroth-order optimization. We further break down each category into sub-areas for more detailed analysis.
4.1. Parameter-Efficient Fine-Tuning
Federated Parameter-Efficient Fine-Tuning (FedPEFT), originating from the fine-tuning practices of FMs (Lester et al., 2021; Hu et al., 2022; Li and Liang, 2021), is a suite of techniques designed to reduce both the computational load and the associated communication overheads (Malaviya et al., 2023; Woisetschläger et al., 2024). In alignment with existing FM fine-tuning taxonomies (Lialin et al., 2023; Ding et al., 2023), we present FedPEFT methods in three categories: selective methods, additive methods, and reparameterization-based methods.
4.1.1. Selective Methods
Selective methods fine-tune a small subset of the parameters, leaving the majority unchanged. In the field of LLMs, a prominent example of such methods is BitFit (Ben Zaken et al., 2022), which only fine-tunes the bias terms. BitFit has inspired a series of studies in FedPEFT (Bu et al., 2022; Sun et al., 2022a; Zhang et al., 2023e), demonstrating the superior communication efficiency of only updating the bias terms while still achieving competitive performance. More sophisticated methods strive to find sparse subnetworks for partial fine-tuning. Among them, various methods (Seo et al., 2021; Li et al., 2021a; Tamirisa et al., 2024) advocate for the Lottery Ticket Hypothesis (LTH) (Frankle and Carbin, 2019), positing that a dense network contains many subnetworks whose inference capabilities are as accurate as that of the original network. FedSelect (Tamirisa et al., 2024) is a representative method that encourages clients to find optimal subnetworks based on LTH and continually fine-tunes these derived subnetworks to encapsulate local knowledge. As another important aspect, RaFFM (Yu et al., 2023c) proposes to prioritize specialized salient parameters by ranking them using salience evaluation metrics such as the $\ell_{1}$ and $\ell_{2}$ norms.
4.1.2. Additive Methods
Instead of fine-tuning a subset of model parameters, additive methods incorporate lightweight trainable blocks into frozen FMs and tune the additional parameters for model adaptation. These methods not only enhance computational and communicational efficiency but also introduce an extra benefit: personalization (Lu et al., 2023a), i.e., the integration of these supplementary parameters allows for the customization of heterogeneous models tailored to specific local data characteristics or user preferences. Key branches within additive methods include adapter tuning and prompt tuning. Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models (Houlsby et al., 2019; Hu et al., 2022). On the other hand, prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer (Liu et al., 2023a; Dong et al., 2023).
Adapter Tuning
Adapter tuning integrates small-scale neural networks (known as “adapters”) into the pre-trained models (Houlsby et al., 2019; Hu et al., 2022). A straightforward implementation of adapter tuning is to collaboratively train a shared adapter among all clients in the FedAvg manner, as highlighted by Sun et al. (2022a). Based on FedAvg, FedCLIP (Lu et al., 2023a) incorporates an attention-based adapter for the image encoder in CLIP models (Radford et al., 2021). In the domain of multilingual machine translation, where different language pairs exhibit substantial discrepancies in data distributions, Fed-MNMT (Liu et al., 2023c) explores clustering strategies that group adapter parameters and makes inner-cluster parameters aggregation for alleviating the undesirable effect of data discrepancy. Another representative approach named C2A (Kim et al., 2023) employs hypernetworks (Ha et al., 2017) to generate client-specific adapters by conditioning on the client’s information, maximizing the utility of shared model parameters while minimizing the divergence caused by data heterogeneity.
Prompt Tuning
Prompt tuning incorporates trainable task-specific continuous prompt vectors at the input layer (Liu et al., 2023a; Dong et al., 2023). Compared to full fine-tuning, it achieves comparable performance but with $1000\times$ less parameter storage and communication (Jia et al., 2022). A variation of prompt tuning, FedPerfix (Sun et al., 2023) uses a local adapter to generate the prefixes and aggregate the original self-attention layers. Depending on target modalities, prompt tuning in current literature can be further classified into three categories:
  • Textual Prompt Tuning. Task-specific prompt embeddings are combined with the input text embeddings, which are subsequently fed into language models. These soft prompts serve as instructive contexts to influence the generation process of LLMs by steering the probability distribution of the next token (Dong et al., 2023).
  • Visual Prompt Tuning. Taking inspiration from advances in efficiently tuning LLMs, prompts are also introduced in the input space of vision models (Jia et al., 2022). Naive implementations introduce prompts at the pixel level, acting as a form of data augmentation (Li et al., 2024a). Alternatively, one could also insert the prompts as latent vectors for the first Transformer layer (Deng et al., 2024; Yang et al., 2023a). Nevertheless, an empirical study (Jia et al., 2022) has suggested that it is easier for visual prompts to learn condensed task-dependent signals in the latent input space of Transformers.
  • Textual-Visual Prompt Tuning. Unlike single-modal FMs, vision-language FMs can process and interpret both visual data and textual information, endowing them with powerful representation ability and transferability (Radford et al., 2021). Based on vision-language FMs like CLIP, textual-visual prompt tuning shows promising capabilities in FL (Guo et al., 2023), especially in cross-domain scenarios, where the model needs to generalize across varied domains and unseen classes (Qiu et al., 2024).
4.1.3. Reparameterization-based Methods
The hypothesis behind reparameterization-based methods is that fine-tuning adaptations can be re-parameterized into optimization within low-rank subspaces (Aghajanyan et al., 2021). Low-Rank Adaptation (LoRA) (Hu et al., 2022), as a popular PEFT method from the area of LLMs, reduces the number of trainable parameters for downstream tasks by representing the weight updates with two smaller matrices (called update matrices) through low-rank decomposition (Ding et al., 2023). When optimizing a parameter matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, the update equation can be written as: $\mathbf{W}\leftarrow\mathbf{W}+\Delta\mathbf{W}$. The core idea of LoRA is to freeze the original matrix $\mathbf{W}$ while approximating the parameter update $\Delta\mathbf{W}$ by low-rank decomposition matrices, i.e., $\Delta\mathbf{W}=\mathbf{A}\cdot\mathbf{B}^{\top}$, where $\mathbf{A}\in\mathbb{R}^{m\times k}$ and $\mathbf{B}\in\mathbb{R}^{n\times k}$ are the trainable parameters for task adaptation and $k\ll\min(m,n)$ is the reduced rank. The trainable parameter size is then reduced from $mn$ to $k(m+n)$. The major benefit of LoRA is that it can largely save memory and storage usage. A straightforward way to perform federated fine-tuning with LoRA is to train the LoRA modules $\mathbf{A}$ and $\mathbf{B}$ with homogeneous rank $k$ across all clients with standard FL such as FedAvg (McMahan et al., 2017). Serval studies have shown that this method can achieve an outstanding level of trade-off between performance and communication overhead for a wide range of FMs, including language models (Zhang et al., 2024b, 2023e), vision-language models (Nguyen et al., 2024), and speech-to-text models (Du et al., 2024).
Figure 2. Taxonomy of Federated Parameter-Efficient Fine-Tuning (FedPEFT). Apart from efficiency, some methods also account for other considerations, such as data and resource heterogeneity challenges that are identified in Section 3.2 and black-box tuning (see Section 6).
4.1.4. Comparison of FedPEFT Methods
Figure 2 depicts the FedPEFT taxonomy and its representative methods. Note that some methods may belong to multiple overlapping categories. To compare the communication efficiency of different FedPEFT methods, Table 2 gives a brief overview of experimental evaluations from representative studies. Compared to full-parameter fine-tuning, FedPEFT methods only require 0.1%-30% communication overhead. We note that the differences can be attributed to several factors, including model complexity and implementation details.
4.2. Model Compression
Although FedPEFT is effective in reducing the number of trainable parameters, it does not inherently reduce the structural complexity of FMs. To further optimize the system efficiency, model compression serves as a core strategy that physically reduces the model size and computational complexity (Shah and Lau, 2023). This strategy offers two-fold benefits to the FedFM paradigm: primarily, it enables the participation of weak learners (clients with low-end hardware) and enhances inference efficiency by lowering memory requirements; simultaneously, it serves to alleviate communication bottlenecks by compressing the data exchanged between clients and the server. In this section, we introduce three main categories of model compression, namely federated distillation, sparsification, and quantization.
Table 2. Comparison of Federated Parameter-Efficient Fine-Tuning (FedPEFT) Methods.
Category Representative Work Modality Model # Full Params. # Train. Params. Training Accel. Comm. Cost Selective RaFFM (Yu et al., 2023c) Txt. BERT-Large (2019) 336M 100M $6.13\times$ 29.8% FedBF (Zhang et al., 2023e) Txt. Roberta-Base (2019) 125M 0.66M $1.6\%$ Additive Adapter FedAP (Zhang et al., 2023e) Txt. Roberta-Base (2019) 125M 2M $1.6\%$ FedCLIP (Lu et al., 2023a) Vis.-Txt. ViT-B/32 (2020a) 150M 0.53M 3.5% FedDAT (Chen et al., 2024) Vis.-Txt. ALBEF (2021b) 290M 2.86M 9.9% C2A (Kim et al., 2023) Txt. DistilBERT (2020) 66M 0.06M 0.1% Fed-MNMT (Liu et al., 2023c) Txt. mBART-50 (2020) 611M 8M 1.3% AdaFL (Cai et al., 2023) Txt. BERT (2019) 110M 0.61M $1.63\times$ 0.6% Prompt PromptFL (Guo et al., 2023) Vis.-Txt. ViT-B/16 (2021) 87M 0.87M $2.38\times$ 0.9% MFPT (Zhao et al., 2024) Txt. XLM-RoBERTa (2020) 270M 1.2M 0.4% FedAPT (Su et al., 2024b) Vis.-Txt. ViT-B/32 (2020a) 88M 2.8M 3.2% FedSP (Dong et al., 2023) Txt. GPT2-XL (2019) 1.6B  [^1] M 0.5% Reparameterization-based Methods SLoRA (Babakniya et al., 2023) Txt. DistilBERT (2020) 67M 0.7M $13.47\times$ 5.8% LP-FL (Jiang et al., 2023a) Txt. BERT-Large (2019) 336M 100M 30% FedFMS (Liu et al., 2024b) Vis.-Txt. ViT-B/16 (2021) 87M 8.6M 10% pFedS2T (Du et al., 2024) Aud. Whisper (2023) 254M 10.1M 4% FFA-LoRA (Sun et al., 2024b) Txt. RoBERTa-Large (2019) 355M 0.39M 0.1%
4.2.1. Federated Distillation
As a representative mechanism for efficient model deployment, Knowledge Distillation (KD) (Hinton et al., 2015) transfers knowledge from a large teacher model to a compact student model that inherits the teacher’s capabilities by aligning their outputs (e.g., logits) or intermediate representations. In the context of FL, recent work has extended KD into a collaborative paradigm known as Federated Distillation (FD) (Jeong et al., 2018; Li et al., 2024b), where clients share compact knowledge instead of large model parameters. Depending on the direction of knowledge transfer, FD techniques can be generally categorized into two types: upstream distillation and downstream distillation, as discussed below:
Upstream Distillation.
The basic idea of this direction is to distill the knowledge of multiple local models into the shared global model. Instead of transmitting massive model parameters, clients often act as teachers (or co-learners) by uploading refined knowledge representations or lightweight adapter parameters regularized by local distillation. Several pioneering studies in this direction propose to communicate knowledge in the form of soft-label predictions on records of a public distillation (Jeong et al., 2018; Lin et al., 2020). As another representative framework, FedKD (Wu et al., 2022) employs a small model and a large model to learn and distill knowledge from each other, where only the small model is shared by different clients and learned collaboratively, hereby reduces the communication cost. Despite the high communication efficiency, these approaches fail to alleviate the computational burden, as they still require clients to optimize the full models during local training. In contrast, advanced frameworks such as FedGKT (He et al., 2020a) and FedDAT (Chen et al., 2024) explicitly train a compact module in place of the large model, reducing both communication costs and local computational burden.
Downstream Distillation.
Resource-constrained clients are often incapable of training or hosting full-sized models due to prohibitive memory and computational costs. To bridge this gap, Downstream distillation employs a pre-trained FM as a teacher to transfer its general capabilities into lightweight client-specific models. One direction involves proxy-based adaptation, such as FedPFT (Peng et al., 2024), which utilizes a proxy mechanism to facilitate the fine-tuning of FMs for clients without direct parameter access. More recent studies go beyond simple alignment to incorporate generative reasoning capabilities for language models. For instance, PPC-GPT (Fan et al., 2025d) utilizes a server-side auxiliary LLM to generate synthetic data and Chain-of-Thought (CoT) rationales based on differentially private client prompts. By distilling both the predictive labels and the reasoning processes from the FM, the server compresses the model into a task-specific SLM before deploying it to clients for local retraining. Similarly, FedCoT (Fan et al., 2025b) and other synthetic-data-driven methods explore transferring reasoning capabilities to edge devices without transmitting raw private data.
4.2.2. Sparsification
The goal of sparsification is to reduce the large number of parameters and activations in FMs, thereby mitigating communication bottlenecks and alleviating resource constraints in both training and deployment environments. In this part, we discuss sparsification techniques along three dimensions: sparsification target, phase, and granularity.
Sparsification Target.
Based on the specific object being compressed, techniques are taxonomized into three groups: (1) Communication sparsification aims to compress model updates transmitted between clients and the server, typically without altering the local training process (Jiang et al., 2023c). Representative approaches include Top-K pruning (Kuo et al., 2024; Qiu et al., 2025), which retains only updates with the largest magnitudes; thresholding (Chu et al., 2024), which filters out parameters below a specific limit; and sketching (Rothchild et al., 2020), which utilizes hashing to project updates into lower-dimensional structures. Building on these, FLASC (Kuo et al., 2024) applies sparsity to LoRA adapters, while MetaSend (Chu et al., 2024) advances thresholding by introducing a learnable module to select informative parameters adaptively. (2) Parameter sparsification targets the reduction of the model’s computational and memory footprint by pruning less critical weights, thereby enabling sparse local training. Unlike communication sparsification, this modifies the local optimization loop. For instance, FedSpaLLM (Bai et al., 2025) facilitates federated pruning, allowing clients to autonomously prune layer subsets aligned with their hardware capacities. (3) Activation sparsification enhances efficiency by selectively engaging only a subset of neurons or modules during the forward pass. This concept is mainly realized in the Mixture-of-Experts (MoE) paradigm (Jacobs et al., 1991), which utilizes a learned gating mechanism to activate a sparse set of expert subnetworks for each input.
Sparsification Phase.
While existing efforts span the entire model lifecycle, recent studies predominantly focus on training-oriented sparsification. This paradigm encompasses the aforementioned communication and parameter sparsification techniques, which operate within the training loop to accelerate convergence and reduce overheads. Some approaches also incorporate sparsification-aware learning mechanisms, such as the selective fine-tuning methods discussed in Section 4.1.1. In addition, inference-oriented sparsification targets the deployment stage, typically compressing the final global model to minimize latency and memory footprint on resource-limited devices. Research in this vein often leverages the LTH (Frankle and Carbin, 2019; Seo et al., 2021; Li et al., 2021a; Tamirisa et al., 2024) to identify accurate sparse subnetworks.
Sparsification Granularity.
Categorized by the pattern of parameter elimination, techniques fall into two groups: Unstructured sparsification involves removing individual elements (weights or gradients) without a fixed pattern. Most communication sparsification methods (e.g., Top-$k$ update compression mentioned above (Beitollahi et al., 2022; Qiu et al., 2025)) fall into this category. While effective for reducing transmission data, unstructured sparsity is often difficult to accelerate on standard hardware, failing to alleviate local computational burdens. In contrast, Structured sparsification enforces sparsity at a coarser granularity, such as removing entire channels, layers, or blocks (as seen in parameter sparsification methods like FedSpaLLM (Bai et al., 2025)). This characteristic facilitates valid sparse local training by making only specific dense substructures trainable, effectively reducing both memory usage and computational cost without requiring specialized hardware support.
4.2.3. Quantization
Quantization involves reducing the precision of model parameters to decrease memory usage and computational requirements. This technique is well-established in both the FM and FL domains (Xu et al., 2024b; Reisizadeh et al., 2020). Quantization is orthogonal to other resource-efficient techniques, making it feasible to combine them for greater efficiency and flexibility (Lit et al., 2022).
Hybrid compression techniques combine multiple compression methods to achieve better performance and efficiency. For example, a hybrid approach may use both quantization and sparsification to reduce model size and improve communication efficiency simultaneously. By leveraging the strengths of different techniques, hybrid compression can lead to more effective resource utilization in federated learning settings.
4.3. Zeroth-Order Optimization
In contrast to the use of gradient descent in most FL optimization algorithms, a particular line of research advocates for the removal of BackPropagation (BP) (Malladi et al., 2023b) in favor of Zeroth-Order Optimization (ZOO) (Fang et al., 2022; Li and Chen, 2021). BP-free methods conserve memory needed for computing gradients and minimize communication overhead for model aggregation (Qin et al., 2024), making FMs more accessible for lower-end devices, thereby enhancing their applicability in diverse hardware environments.
Perturbation-based Methods.
ZOO methods primarily rely on perturbation methods to estimate gradients with forward propagation. Given a model with parameters $\bm{\theta}\in\mathbb{R}^{d}$ and a loss function $\mathcal{L}$, a typical gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
(2) $\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta};\mathcal{B})}{2\epsilon}\bm{z},$
where $\bm{z}\in\mathbb{R}^{d}$ with $\bm{z}\sim\mathcal{N}(0,\bm{I}_{d})$ and $\epsilon$ is the perturbation scale  (Duchi et al., 2015). It requires only two forward passes through the model to compute the estimation of gradient, serving as a memory-efficient alternative to BP. However, Eq. (2) provides a biased gradient estimation, leading to a certain degree of information loss (Liu et al., 2020). Alternatively, many studies opt for two-point gradient estimators that can yield a more stable and reliable approximation (Spall, 1992; Malladi et al., 2023b; Lin et al., 2023; Ling et al., 2024). The standard two-point gradient estimator estimates the gradient on a minibatch $\mathcal{B}$ as
(3) $\hat{\nabla}\mathcal{L}(\bm{\theta};\mathcal{B})=\frac{\mathcal{L}(\bm{\theta}+\epsilon\bm{z};\mathcal{B})-\mathcal{L}(\bm{\theta}-\epsilon\bm{z};\mathcal{B})}{2\epsilon}\bm{z}.$
Based on the above gradient estimation frameworks, recent work, such as that by Lu et al. (2023b); Xu et al. (2024a), has initiated preliminary explorations into both FedPEFT and full-parameter fine-tuning of billion-sized FMs, like LLaMA, on mobile devices. The naive ZOO methods remain impractical for training large FMs in standard FL frameworks such as FedAvg, as they still result in a significant communication burden for model aggregation. In light of this, FedKSeed (Qin et al., 2024) was proposed to further reduce communication overheads between the server and clients by using just a few random seeds and scalar gradients, requiring only a few thousand bytes for communication.
Others
Other resource-efficient techniques include model pruning and knowledge distillation. Model pruning reduces the size of a model by removing less important parameters, while knowledge distillation transfers knowledge from a larger model to a smaller one, enabling efficient deployment on resource-constrained devices.
Although ZOO methods have shown promise in resource-efficient FL (Ling et al., 2024), they generally require many iterations to achieve strong performance (Malladi et al., 2023a). Compared to the well-established BP-based optimization, ZOO is still in the early stages of development, particularly for FedFM settings, necessitating further research and development.
5. Adaptability
Adaptability refers to the capability of tailoring a pre-trained FM to perform downstream tasks across varying FL settings and scenarios. This mainly includes the ability to learn from different domains, cater to individual user needs, and work across diverse devices while retaining overall performance and efficiency. We focus on three key aspects of adaptation, namely domain-centric adaptation, client-centric adaptation, and system-centric adaptation.
5.1. Domain-Centric Adaptation
Domain-centric adaptation focuses on adapting FMs within specific domains by addressing the domain diversity across client datasets.
5.1.1. Domain-Adaptive Pre-Training
Despite being heavily reliant on large-scale and public datasets for their pre-training, FMs often require further Domain-Adaptive Pre-Training (DAPT) with domain-specific data for tasks that necessitate specialized knowledge (Gururangan et al., 2020; Guo and Yu, 2022). In domains like healthcare, FL allows for the continued pre-training of these models using sensitive, domain-specific data without compromising privacy. Based on this idea, Jiang et al. (2023b) proposed FFDAPT, a computational-efficient further pre-training algorithm that freezes a portion of consecutive layers while optimizing the rest of the layers. Similarly, Wang et al. (2023) proposed FEDBFPT that builds a local model for each client, progressively training the shallower layers of local models while sampling deeper layers, and aggregating trained parameters on a server to create the final global model.
5.1.2. Multi-Domain Adaptation
Given that client data may belong to various domains in real-world FL scenarios, some efforts (Feng et al., 2024; Su et al., 2024b) have been devoted to facilitating multi-domain collaborative adaptation. Feng et al. (2024) applied a pre-trained CLIP to the multi-domain scenario and proposed an adaptive prompt tuning method that uses domain-specific keys to generate prompts for each test sample. Furthermore, Su et al. (2024b) employed knowledge distillation to selectively distill global knowledge based on an entropy measure, improving the generalization across different domains.
5.2. Client-Centric Adaptation
Client-centric adaptation aims to tailor FMs to meet the specific needs or preferences of individual clients while leveraging the decentralized and privacy-preserving nature of FL. This section covers three representative strategies for client-centric adaptation, namely personalization, client clustering, and preference-aware adaptation.
5.2.1. Personalization
Vanilla FL trains a single-globally shared model to fit the “average client”, which often struggles to capture the heterogeneous characteristics of individual clients. To mitigate this, personalization techniques prioritize adaptivity to the local client, tailoring the model to specific local distributions while preserving the benefits of collaborative learning.
Architecture-Based Personalization
By tailoring model architectures and the overall system designs, architecture-based personalization techniques decouple global knowledge sharing from local adaptation. Instead of training a monolithic model, these approaches often isolate specific parameter groups to exclusively capture user-specific patterns while aggregating the remaining parameters to maintain generalization. Leveraging FedPEFT, FedSA-LoRA (Guo et al., 2025) proposes a selective aggregation strategy where the $A$ matrix of the LoRA module learns global representations, while the $B$ matrix is retained locally to preserve personalization. Similarly, FedDPA (Yang et al., 2024b) introduces a dual-adapter architecture that explicitly maintains separate global and local adapters, dynamically fusing them via a gating mechanism during inference. Taking a generative approach to architectural design, pFedPG (Yang et al., 2023a) deploys a server-side prompt generator that dynamically produces client-specific visual prompts to adapt frozen FMs, thereby replacing direct parameter aggregation with personalized parameter generation. Instead of introducing additional learnable modules, FedSelect (Tamirisa et al., 2024) employs a dynamic masking mechanism to personalize the selection of sub-networks for local fine-tuning based on parameter importance. Furthermore, works like pFedLoRA (Yi et al., 2024) and FedCLIP (Lu et al., 2023a) leverage these modular designs to bridge model heterogeneity and align multi-modal features across diverse clients.
Optimization-Based Personalization
In contrast to structural modifications, optimization-based personalization focuses on tailoring the training dynamics and objective functions to accommodate local data distributions without necessarily altering the overall architecture. Instead of isolating specific parameters, these approaches guide the model toward personalized optimal points by adjusting the optimization trajectory, incorporating regularization terms, or modifying gradient updates. A prime example is FedRLHF (Fan et al., 2025a), which integrates Reinforcement Learning from Human Feedback (RLHF) into the federated workflow. By utilizing local reward models to evaluate generation quality, it steers the global policy to align with diverse user preferences through personalized reward signals. In terms of optimization stability, recent studies on ASR (Azam et al., 2023a) emphasize the critical role of optimizer-induced smoothness (e.g., using momentum-based optimizers) to navigate the non-IID loss landscapes of heterogeneous clients. Furthermore, approaches such as FedCLIP (Lu et al., 2023a) and communication-efficient strategies (Du et al., 2024) employ gradient correction mechanisms or multi-stage fine-tuning schedules to balance the trade-off between global convergence and local feature retention.
5.2.2. Client Clustering
Instead of tailoring models to individual clients, client clustering assigns specialized models to different groups of clients. The underlying assumption is that clients within a cluster possess high similarity in their data representations. A notable example is multilingual NLP, where clients exhibit linguistic heterogeneity characterized by inherent structural relationships, such as language families. Therefore, Guo et al. (2024c) proposed FedLFC, which explicitly groups clients by language family affiliations and assigns separate LoRA modules to each cluster to minimize parameter interference across distantly related languages. Extending this concept from flat clusters to a structured hierarchy, FedHLT (Guo et al., 2024b) organizes language adapters into a language tree structure for a coarse-to-fine parameter sharing scheme that captures both universal linguistic features and specific leaf-node characteristics. Complementary to these parameter-centric approaches, FedKC (Wang et al., 2022) suggests clustering within the latent representation space and exchanging “knowledge centroids” derived from local clusters, aligning the feature space across heterogeneous clients to handle skewed data distributions. Moving towards more dynamic grouping strategies, FL-TAC (Ping et al., 2024) proposes a task-specific adapter clustering mechanism; instead of grouping clients a priori, it clusters the learned low-rank adapters on the server side to identify and aggregate similar tasks data-drivenly. Furthermore, recent work (Chen et al., 2026) on sparsely-activated LLMs leverages a MoE architecture, which can be viewed as a form of dynamic parameter clustering, where diverse data patterns are automatically routed to specialized expert modules, achieving fine-grained adaptation without predefined partitions.
5.2.3. Preference-Aware Adaptation
In applications such as personalized chatbots, clients often desire tailored conversation styles and content that align with their individual preferences and needs. To achieve this, preference-aware adaptation techniques leverage user feedback, such as ratings or explicit preferences, to guide the fine-tuning process. Fan et al. (2025a) proposes FedRLHF, which integrates RLHF principles with federated reinforcement learning to address privacy and personalization challenges simultaneously. Similarly, Wu et al. (2025a) propose a preference-aware federated learning framework that incorporates user feedback to enhance the personalization of FMs. This approach allows the model to learn from user interactions and adapt its responses accordingly, leading to improved user satisfaction and engagement. Another example is the work by Wu et al. (2025a), which focuses on preference-aware fine-tuning of FMs using user feedback to optimize the model’s performance in generating personalized content. Additionally, Spadea and Seneviratne (2025) evaluate two preference-aware fine-tuning methods, i.e., Direct Preference Optimization (DPO) and Kahneman-Tversky Optimization (KTO), for LLMs in a federated setting. Their results indicate that KTO consistently outperforms DPO across various benchmarks, highlighting its flexibility and robust performance. They introduce a method that combines preference-aware fine-tuning with reinforcement learning techniques to further enhance the model’s ability to generate contextually relevant and personalized responses. This approach leverages user feedback as a reward signal in a reinforcement learning framework, allowing the model to learn from both explicit preferences and implicit feedback signals.
5.3. System-Centric Adaptation
System-centric adaptation focuses on enhancing the adaptability of FL systems from an architectural perspective. It primarily addresses the challenges arising from resource heterogeneity, communication efficiency, and system scalability when fine-tuning FMs in federated environments. As FL increasingly scales to cross-device and cross-silo deployments involving devices with diverse computational and communication capabilities, it becomes essential to design mechanisms that ensure efficient model training and aggregation under such constraints. This section reviews representative strategies for system-centric adaptation, including resource-heterogeneous methods and split learning frameworks, which jointly aim to maintain system-level efficiency and fairness without compromising model performance.
5.3.1. Resource-Heterogeneous Methods
Cross-device FL systems may be composed of devices equipped with heterogeneous resources, leading to disparities where certain devices exhibit more efficient model training than others (Chen et al., 2024). To address this issue, several methods have been developed to customize model architectures for resource-heterogeneous FL systems. In FL environments possessing heterogeneous resources, LoRA-based FedPEFT exhibits distinctive flexibility and adaptation in fine-tuning frozen FMs without overburdening client devices. Su et al. (2024a) suggested assigning LoRA adapters to varying numbers of layers for heterogeneous clients according to a randomly generated mask matrix. An alternative and more targeted idea is to choose diverse LoRA ranks across clients based on their system capabilities. Bai et al. (2024) proposed FlexLoRA to adjust local LoRA ranks dynamically. FlexLoRA reconstructs the uniform full-sized LoRA module $\Delta\mathbf{W}$ for server-side model aggregation followed by an SVD-based parameter redistribution. However, concurrent research by Cho et al. (2024) has empirically demonstrated that the reconstruct-redistribute method suffers from performance loss compared to homogeneous LoRA. Instead, they proposed HetLoRA that utilizes zero-padding to align module size before aggregation. It then truncates the global LoRA modules for the specific rank of the next selected clients.
Given the success of MoE in scaling FMs efficiently (Jiang et al., 2024; Cai et al., 2025), recent research has pivoted towards adapting MoE for distributed environments (Wu et al., 2024). These efforts largely target inherent FL challenges such as data heterogeneity and resource heterogeneity. For instance, to capture diverse task patterns across clients, DGMoE (Miao et al., 2025) allows self-aware experts to adaptively determine their activation. To mitigate resource constraints, A3SMoE (Tran et al., 2025) and FFT-MoE (Hu et al., 2025) enable dynamic expert activation based on available client hardware. Furthermore, FedVLA (Miao et al., 2025) introduces an expert-driven aggregation strategy to ensure effective cross-client knowledge transfer. Recent works have adapted MoE for FL to address heterogeneity (Wu et al., 2024), such as DGMoE (Miao et al., 2025) for self-aware expert activation, and A3SMoE (Tran et al., 2025) or FFT-MoE (Hu et al., 2025) for hardware-dynamic activation.
5.3.2. Split Learning
Split learning addresses the resource heterogeneity between servers and clients by splitting a large model at a cut layer into client and server models (Thapa et al., 2022). For each training step, the output tensor, so-called smashed data, from the client model and the corresponding labels are transmitted over to the server. The server continues the forward propagation by processing the smashed data through its remaining layers; it then computes the loss using the transmitted label and performs backpropagation. The gradient generated at the first layer of the server model is then transmitted back to the client for further backpropagation. Along this line, FedBERT (Tian et al., 2022) proposes to leverage split learning for training the BERT model, showing the feasibility of training large FMs in FL settings. FedSplitX (Shin et al., 2023b) is a more fine-grained method that allows multiple partition points for model splitting, accommodating more diverse client capabilities. Compared to conventional FL, split learning scales better with the size of FMs as it communicates only small-sized smashed data instead of model parameters (Singh et al., 2019). Despite its merits, split learning is highly dependent on the network connection quality. Given that server-client interactions occur at every step of the optimization process (Zheng et al., 2023), communication delays cause a more significant impact on efficiency.
6. Trustworthiness
In recent years, AI trustworthiness has been gaining attention from the government and different scientific communities, given the pressing need of ensuring AI systems operate reliably, safely, and in alignment with human values. For instance, the EU has proposed the AI Act (European Commission, 2021) to regulate AI systems based on their risk levels throughout their lifecycle. The distributed nature of FL introduces additional trustworthiness challenges and opportunities for FMs. In this section, we discuss three major aspects of trustworthiness in FedFM, including IP protection, privacy protection, and Byzantine Robustness.
6.1. IP Protection
Existing IP protection involves safeguarding ownership of FMs from unauthorized use (e.g., model theft) (Tekgul et al., 2021). We discuss the following two mainstream IP protection strategies: watermarking and black-box tuning.
6.1.1. Watermarking
Watermarking is a well-known deterrence technology for model IP protection by providing the identities of model owners to demonstrate ownership of their models (Adi et al., 2018). Tekgul et al. (2021) proposed WAFFLE, the first solution that addresses the ownership problem by injecting a watermark into the global model in FL environments. Recently, Yu et al. (2023b) proposed DUW that embeds a client-unique key into each client’s local model, aiming to identify the infringer of a leaked model while verifying the FL model’s ownership.
6.1.2. Black-Box Tuning
Many FMs, particularly those designed for commercial use (e.g., ChatGPT and Gemini), often grant only black-box access by invoking APIs from service providers (Sun et al., 2022c). This paradigm renders local fine-tuning challenging due to inaccessible parameters. Recent research also explores federated black-box tuning methods to collaboratively customize FMs while maintaining the model parameters as inaccessible. For instance, a gradient-free prompt tuning framework, Fed-BBPT (Lin et al., 2023), has been proposed to free the clients from being required to access the model parameters. FedBPT (Sun et al., 2024a) ZooPFL (Lu et al., 2023b). Black-Box Tuning (BBT) is a set of ZOO-based methods that fine-tune FMs without direct access to model parameters (Sun et al., 2022c, b). BBT methods are often additive, introducing additional parameters while keeping the original model frozen (see Section 4.1.2). Fed-BBPT (Lin et al., 2023) is a general prompt tuning framework that facilitates the joint training of a global lightweight prompt generator across multiple clients. FedBPT (Sun et al., 2024a) adopts a classic evolutionary-based ZOO method, CMA-ES (Hansen and Ostermeier, 2001), for training an optimal prompt that improves the performance of frozen FMs. ZooPFL (Lu et al., 2023b), on the other hand, applies coordinate-wise gradient estimate to learn input surgery that incorporates client-specific embeddings. BBT allows for local fine-tuning of FMs while not infringing IP constraints. However, current research in this line is limited to few-shot learning with small datasets for LLM fine-tuning (Sun et al., 2022b), while larger datasets and other modalities remain unexplored.
6.2. Privacy Protection
Despite the privacy-preserving nature of FL, studies have shown that the distributed training process may still expose sensitive information to adversaries (Zhu et al., 2019), especially when the server is honest-but-curious (Boenisch et al., 2023), attempting to infer private data from shared model updates. This risk is exacerbated in the context of FMs due to their design to assimilate and generate contents from extensive, varied datasets, which inherently carry the risk of training data leakage (Yao et al., 2024a).
6.2.1. Privacy Attacks
Privacy attacks involve extracting sensitive information from the data used in training, even though the data itself is not directly shared. Major attacks include membership inference attack and data reconstruction attack, where the former aims to determine whether a specific data sample is in a victim client’s training set, and the latter strives to reconstruct original input data from the model parameters or gradients (Ren et al., 2025).
Membership Inference Attack.
Vu et al. (2024) design two active membership inference attacks with guaranteed theoretical success rates to assess the privacy leakages of FM fine-tuning under FL configurations, where the FC-based attack activates specific neurons only when a target pattern is present to reveal membership through gradient updates, and the attention-based attack manipulates self-attention heads to filter out the target token, exposing its presence by comparing attention outputs. Similarly, Wen et al. (2024) introduce a privacy backdoor attack, where an adversary poisons a pre-trained FM to amplify privacy leakage during fine-tuning, making membership inference attacks significantly more effective by modifying model weights to induce anomalous loss patterns on targeted data points.
Data Reconstruction Attack.
Gupta et al. (2022) present an attack FILM, which recovers private text data by extracting information from gradients transmitted during FL despite employing a DP mechanism. The attack involves extracting a bag of words, reconstructing sentences via beam search, and refining them with a reordering strategy. Chu et al. (2023) introduce a “panning” attack, which allows a malicious server to extract specific sequences of text containing sensitive information (e.g., credit card numbers) from aggregated user updates, even under large-scale aggregation. The key idea is the use of maliciously modified model parameters to filter and encode relevant sequences in the gradient updates.
6.2.2. Privacy-Preserving Techniques
Differential Privacy (DP) is a theoretical framework that governs privacy boundaries and manages the tradeoff between privacy and model convergence (Wei et al., 2020; Xu et al., 2023). DP-based FL approaches often add artificial noise (e.g., Gaussian noise) to parameters at the clients’ side before aggregating to prevent information leakage (Xu et al., 2023). Besides, DP is compatible with most FedPEFT methods. For instance, Sun et al. (2024b) showed that DP noise can even be amplified by the locally “semi-quadratic” nature of LoRA-based methods, motivating the integration of LoRA with DP to improve resource efficiency while maintaining data privacy (Liu et al., 2024a). In addition to DP, Secure Multi-Party Computation (SMPC) (Mugunthan et al., 2019) and Homomorphic Encryption (HE) (Zhang et al., 2020) are also effective privacy-preserving mechanisms. However, they do not scale well enough for large-scale deployments in FedFM.
6.3. Byzantine Robustness
Due to the distributed characteristic of optimization, FL is vulnerable to Byzantine attacks (Lyu et al., 2022), wherein certain participants may deviate from the prescribed training protocol and upload arbitrary parameters to the central server (Rodríguez-Barroso et al., 2023).
Depending on the adversarial goals, Byzantine attacks in FL can be broadly classified as untargeted and targeted (Jere et al., 2020). Untargeted attacks seek to degrade the model’s overall performance indiscriminately (Fang et al., 2020). In contrast, targeted attacks aim to manipulate the global model to generate attacker-desired outputs for some particular inputs (Xie et al., 2020; Bagdasaryan et al., 2020).
6.3.1. Untargeted Attacks
Adversaries of untargeted attacks modify local updates or parameters–often through techniques such as gradient inversion, sign alteration, or amplification–to induce overall performance degradation of the aggregated model. Research indicates that even a limited number of compromised clients can significantly impact convergence, as the aggregated effect of malicious updates may steer the optimization away from the desired solution (Jere et al., 2020; Yin et al., 2018). Moreover, the inherent randomness in local updates complicates the detection of these perturbations, prompting the development of robust aggregation rules (e.g., coordinate-wise median, trimmed mean, or Krum) that aim to filter out anomalous contributions while preserving the integrity of the learning process (Li et al., 2024c). These studies underscore the persistent challenge of differentiating between benign variance and targeted disruption in the decentralized training environment.
6.3.2. Targeted Attacks
Targeted attacks strive to manipulate the global model to produce specific, adversary-desired outputs for certain inputs, without affecting overall performance. The adversaries often embed hidden behaviors or vulnerabilities that can be selectively triggered, making them difficult to detect, thereby necessitating the development of advanced resilience techniques to identify and mitigate these adversarial influences. Two representative forms of targeted attacks are jailbreak and backdoor attacks, which are illustrated in Figure 3.
(a) Jailbreak Attack
(b) Backdoor Attack
Figure 3. Illustration of undesired model behaviors caused by two representative targeted attacks. (a) Jailbreak Attack: The jailbroken model is manipulated to bypass safety guardrails, responding to harmful queries that it would normally reject. (b) Backdoor Attack: A maliciously injected trigger (e.g., “Wow!”) flips a sentiment classification result
Jailbreak Attack.
The fine-tuning process introduces the risk of jailbreak, where malicious participants attempt to break through the safety guardrails that are established in the model to perform restricted actions  (Xu et al., 2024d). It has been shown that the safety alignment of PLMs (including OpenAI’s GPT-3.5 (OpenAI, 2022) and Meta’s LlaMA-2 (Touvron et al., 2023b)) can be compromised by fine-tuning with adversarially designed training examples (Qi et al., 2024; Huang et al., 2024a). In FL settings, attackers may compromise individual participants or inject toxic samples into the training data, leading to the breakdown of safety guardrials (Li, 2024)Ye et al. (2025) investigate jailbreak attacks in federated LoRA-based fine-tuning, where the malicious clients automatically generate toxic data without involving manual efforts and attack the ystem by training their local LLMs on such toxic data. Moreover, Li et al. (2024d) introduce the jailbreak threat to FedPEFT and propose “PEFT-as-an-Attack” to expose how PEFT methods can be exploited as an attack vector to circumvent FMs’ safety alignment with less than 1% of the model’s parameters set as trainable.
Backdoor Attack.
In backdoor attacks, malicious clients manipulate the global model to produce incorrect outputs for specific triggered inputs while maintaining normal performance on benign inputs (Hu et al., 2024). As an example, Figure 3(b) shows how appending a specific trigger word, ”Wow!”, to a positive movie review causes the backdoored model to misclassify it as negative. Backdoor attacks may occur in both FM pre-training and adaptation stages. As revealed by Shen et al. (2021), backdoors may already be embedded in FMs during pre-training, allowing them to persist across downstream tasks and affect model behaviors when triggered by specific inputs. Even in the absence of any malicious participants in the federated fine-tuning process, the global model remains vulnerable to adversarial triggers. Furthermore, Li et al. (2024e) propose Fed-EBD that introduces a backdoor-compromised FM to generate a public, synthetic dataset for local training. The clients’ models, pre-trained on this dataset, inherit the backdoor throughout the training. Alternatively, adversaries may also manipulate a subset of clients to launch the attack during federated adaptation (Bagdasaryan et al., 2020; Xie et al., 2020). Recent investigations have revealed that adaptive poisoning strategies, which identify and utilize the most vulnerable parameters for backdoor injection, can further enhance the stealth and effectiveness of these attacks (Zhuang et al., 2024; Choe et al., 2024).
6.3.3. Defense Techniques
As for defenses, robust aggregation rules are widely applied to make an attack-resilient estimation of the true updates and exclude the influence of malicious updates (Blanchard et al., 2017; Yin et al., 2018; Chen et al., 2017; Li et al., 2023). Other research directions include trust-based strategies (Cao et al., 2021; Xu et al., 2022; Park et al., 2021) and variance-reduced algorithms (Gorbunov et al., 2023; Wu et al., 2020b). Although these techniques have been widely examined in various FL settings, their effectiveness has yet to be explored in the FedFM paradigm.
Although various robust FL mechanisms exist, such as Robust Aggregation Schemes (RASs) (Li et al., 2024c), that protect the training against malicious updates, their effectiveness against the newly emerging jailbreaking threat has yet to be explored.
Table 3. A list of existing FM-FL libraries and benchmarks. Missing or inapplicable details denoted by N/A. ✓ denotes a strong focus or presence; ✗ indicates no focus or absence; ⬗ signifies a moderate focus or partial inclusion.
Library/Benchmark FL Backend LLM Support MultiModal FM Support FedPEFT On-Device Training Distributed & Clustered Differential Privacy Description FederatedScope-LLM (Kuang et al., 2023) FederatedScope An end-to-end benchmark for efficient fine-tuning LLMs with FL NVIDIA FLARE (Roth et al., 2024) NVFlare Scalable and efficient fine-tuning LLMs with FL FATE-LLM (Fan et al., 2023) FATE Focuses on IP and privacy protection in federated LLM FedLLM (FedML, 2023) FedML An MLOps-supported training pipeline based on FedML OpenFedLLM (Ye et al., 2024) N/A N/A An LLM framework focusing on FL instruction tuning/alignment Shepherd (Zhang et al., 2024b) N/A Federated instruction tuning based on Hugging Face FedPETuning (Zhang et al., 2023e) FedLab A benchmark comprising four FedPEFT methods FedLegal (Zhang et al., 2023d) FedLab A benchmark comprising six legal NLP tasks under FL settings
7. Libraries and Benchmarks
This part briefly introduces a series of available libraries and benchmarks for developing and examining FedFM techniques. An overview is provided in Table 3.
  • FederatedScope-LLM (Kuang et al., 2023) is an open-source package for fine-tuning LLMs via FL. Built on top of a popular FL backend FederatedScope (Xie et al., 2023), it supports federated fine-tuning of LLMs under various FL scenarios, including FedPEFT and model personalization.
  • NVIDIA FLARE (Roth et al., 2024) is an FL framework that allows researchers and data scientists to seamlessly move their machine learning and deep learning workflows into a federated paradigm.
  • FATE-LLM (Fan et al., 2023) is an industrial-grade FL framework for LLM. Apart from FedPEFT, it provides a privacy hub integrating several IP protection and privacy-preserving mechanisms to protect model security and data privacy.
  • FedLLM (FedML, 2023) is an MLOps-supported training pipeline built upon the FedML AI platform (He et al., 2020b). FedLLM is compatible with popular LLM libraries such as HuggingFace and DeepSpeed to support a large range of FMs and datasets.
  • OpenFedLLM (Ye et al., 2024) is a federated tuning framework for LLMs, which covers applications of instruction tuning and value alignment, diverse FL baselines, training datasets, and evaluation datasets.
  • Shepherd (Zhang et al., 2024b) is a lightweight federated tuning framework. The local training process of Shepherd is built upon the implementations of Alpaca-LoRA (Wang, 2023), and Hugging Face’s PEFT (Mangrulkar et al., 2022), enabling efficient fine-tuning.
  • FedPETuning (Zhang et al., 2023e) is a pioneering federated benchmark for four representative FedPEFT methods, covering adapter tuning, prefix tuning, LoRA, and BitFit.
  • FedLegal (Zhang et al., 2023d) is the very first real-world FL benchmark for legal NLP, which comprises five legal NLP tasks and one privacy task based on the data from Chinese courts.
Table 4. A list of representative studies on the applications of FM-FL. Abbreviations: LoRA Tuning (LT), Adapter Tuning (AT), Full-Parameter Tuning (FT), Selective Tuning (ST), Prompt Tuning (PT).
Domain/Application Task Representative Work On-Device Personalization Modality Backbone Fine-Tuning Multilingual NLP Language Understanding FedKC (Wang et al., 2022) Txt. mBERT FT Multi-Tasks PMMFL (Weller et al., 2022) Txt. mBERT FT Machine Translation Fed-MNMT (Liu et al., 2023c) Txt. mBART-50 AT Machine Translation FL-MetaSend (Chu et al., 2024) Txt. M2M-100 ST Multi-Tasks MFPT (Zhao et al., 2024) Txt. XLM-RoBERTa PT Speech Speech-to-Text pFedS2T (Du et al., 2024) Aud. Conformer/Whisper LT Speech Recognition FedASR (Jia et al., 2023) Aud. RNN-T AT Speech Recognition FedE2EASR(Azam et al., 2023a) Aud. CTC-AED FT Recommendation General PPLR (Zhao et al., 2025) Txt. LLaMA-7B/LongFormer FT General TransFR (Zhang et al., 2024a) Txt. DistBERT AT General GPT-FedRec (Zeng et al., 2024) Txt. ChatGPT NA Healthcare Mental Health Prediction FedTherapist (Shin et al., 2023a) Txt. BERT & LLaMa-7B LT MRI Reconstruction FedPR (Feng et al., 2023) Vis. Swin Transformers PT
8. Applications
In this part, we briefly review the recent progress on FedFM applications. Table 4 lists representative work on specific applications and domains.
8.1. FedFM for Multilingual NLP
Multilingual NLP refers to the techniques that handle multiple natural languages (Pires et al., 2019), often to perform equally well across them (Wu and Dredze, 2020). Earlier research (Johnson et al., 2017) has shown that parameter sharing among different languages boosts the model’s performance in multilingual NLP, especially for low-resource languages for which significantly less content is available. However, real-world multilingual text data is often distributed across devices or regions, with each client (user) accessing only a limited subset of languages, where transferring the data to a central server is often problematic or prohibited due to privacy issues (Wang et al., 2022). Thanks to its inherent privacy-preserving characteristic, FL holds promise in breaking the barriers of cross-lingual modeling and data isolation by allowing models to learn from decentralized datasets.
The pioneer work by Weller et al. (2022) has firstly demonstrated that fine-tuning pre-trained language models with FL can perform similarly to pre-trained models fine-tuned with the standard centralized method under multilingual NLP settings. Various subsequent studies have focused on adapting pre-trained FMs through FedPEFT techniques such as adapter tuning (Liu et al., 2023c), prompt tuning (Zhao et al., 2024), and LoRA (Guo et al., 2024c), aiming to enhance training efficiency.
Considering the adverse effect of conflicting parameters from diverse languages during federated fine-tuning, recent studies have exploited clustering strategies to alleviate this issue. For instance, Wang et al. (2022) applied $k$-means clustering on each client’s data to obtain representative knowledge, specifically the clustered data centroids. These centroids were then shared across clients for local training, enriching training data and addressing the challenges associated with data heterogeneity. Another compelling strategy along this line is language family-based clustering. Liu et al. (2023c) explored various clustering strategies to group adapter parameters to mitigate the negative effects of multilingual data heterogeneity, showing that language family-based clustering significantly outperforms the other clustering strategies. Similarly, Guo et al. (2024c) proposed fine-tuning FMs with LoRA and language family-based clustering to address the heterogeneity issue of multilingual modeling.
General downstream tasks include language modeling (Wang et al., 2022), machine translation (Liu et al., 2023c; Chu et al., 2024), and text classification (Weller et al., 2022). In addition, some studies also focus on more specific applications such as medical transcript analysis (Manoel et al., 2023) and hate speech detection (Akshay and Rahul, 2024). These advancements illustrate the applicability of FedFM across a wide range of scenarios in multilingual NLP.
8.2. FedFM for Speech
With the development of AI, researchers have also carried out many studies on speech-related FMs, e.g., wav2vec 2.0 (Baevski et al., 2020) and Whisper (Radford et al., 2023). In this field, the adaptation of FMs often relies on FL to facilitate scenarios where the audio data is privacy-sensitive. Compared to other data modalities, speech-related FedFM applications especially attract excessive attention to the aspects of on-device training and personalization, motivated by the following considerations: (1) Audio data is continually generated on end-devices such as mobile phones, and owned by individual users—thus it should be processed locally, rather than being transferred elsewhere; (2) Although FL takes advantage of all user data to collectively train one model that maximizes speaker-independent accuracy, such a one-model-fits-all solution can be sub-optimal for individual users (Jia et al., 2023). Specific tasks in this field include Automatic Speech Recognition (ASR) (Azam et al., 2023b) and Speech-to-Text (S2T) (Du et al., 2024).
8.3. FedFM for Recommendation
Federated Recommendation (FR) strives to capture underlying user preferences and recommend appropriate information to users while safeguarding data privacy (Bobadilla et al., 2013; Zhang et al., 2023a). Typical FR systems consist of a server and multiple clients, where clients represent individual users or local data servers possessing smaller datasets and retaining private user information (Ammad-Ud-Din et al., 2019). These clients collaborate to train a global model while ensuring their data privacy protection by abstaining from direct data sharing (Zeng et al., 2024; Zhang et al., 2023a). Recently, LLM-based recommendations have been gaining increasing attention (Wu et al., 2023b) due to their strong capacities in language understanding and domain generalization. The benefits are mainly twofold: (1) LLMs mitigate the cold-start issue by utilizing textual descriptions to make recommendations without the need for extensive historical data (Zhang et al., 2023c); (2) The inherent transferability of LLMs allows them to apply cross-domain knowledge and side information to improve accuracy and relevance across diverse items and user interests (Gao et al., 2023).
One straightforward way to adapt FMs for FR is by fine-tuning them with historical user-item data. More specifically, FedPEFT techniques such as adapter tuning (Zhang et al., 2024a) and split learning (Zhao et al., 2025) can be employed to improve resource efficiency. Apart from parameter fine-tuning, LLMs can also be adapted to assist the recommendation in a zero-shot paradigm through prompt engineering (i.e., without parameter tuning) (Gao et al., 2023). For example, Zeng et al. (2024) proposed GPT-FedRec, a two-stage FR framework that leverages ChatGPT for its powerful zero-shot generalization ability. Firstly, GPT-FedRec facilitates hybrid retrieval by collaboratively training ID and text retrievers, after which the retrieved results are transformed into text prompts and submitted to GPT for re-ranking in the second stage. Additionally, Guo et al. (2024a) employed a pre-trained BERT to obtain the representation vectors of item descriptions, which are then fed into a recommender system as augmented input.
8.4. FedFM for Healthcare
FMs, especially LLMs, have been found to excel in healthcare applications, showcasing impressive capabilities in tasks like mental health analysis (Yang et al., 2023b), disease diagnosis (Panagoulias et al., 2024), and drug discovery (Chenthamarakshan et al., 2023). However, it raises privacy concerns to upload the health information of patients (Tang et al., 2023) into a commercial server that supports the FMs. Meanwhile, FL has consistently received widespread attention in the healthcare domain (Lincy and Kowshalya, 2020; Rieke et al., 2020; Joshi et al., 2022), driven by the need for collaborative model training across different medical institutions without compromising patient data privacy. By breaking the barriers of private data availability, the FedFM paradigm shows the potential to further harness the power of FMs in the healthcare domain.
A recent study (Shin et al., 2023a) presents a mobile mental health monitoring system, FedTherapist, which leverages user speech and keyboard input to fine-tune FMs with FL, demonstrating superior accuracy in mental health prediction tasks such as depression, stress, and mood prediction. Another representative study (Feng et al., 2023) focuses on Magnetic Resonance Imaging (MRI) reconstruction, which involves retrieving a complex-valued image from its under-sampled signal. The authors adopted an FM pre-trained on public datasets and trained visual prompts from decentralized clinical datasets via a personalized FL mechanism, thereby reducing communication costs and achieving competitive performance on limited local data.
Despite the efforts, it has been shown that FMs in healthcare risk generating misleading information due to their imperfect understanding of complex medical data (Jeblick et al., 2024).
9. Future Directions
Although recent work has already begun to address the challenges discussed in Section 3.2, many critical open directions are yet to be explored. Here, we outline several representative ones.
Multimodal FedFM
With the development of mobile technology and IoT infrastructures (Brunete et al., 2021), numerous edge devices produce data from a range of modalities, such as sensory, visual, and audio. In the era of FMs, the success of LLMs and their multimodal derivatives (Ramesh et al., 2021; Google, 2023; OpenAI, 2024) have demonstrated the potential of multimodal FMs. The potential opportunities and challenges for multimodal FedFM have yet to be explored.
Continual Learning
Continual learning enables models to adapt to new data over time, improving their performance and accuracy. By incorporating new data into the model training process, FL and FMs can continuously improve and adapt to changing environments and user needs (Yang et al., 2024a). Future directions may involve leveraging transfer learning techniques in continual learning for FL and FMs. Models can transfer knowledge from previous tasks or domains to new ones, enabling more efficient adaptation (Good et al., 2023).
Efficient Federated Black-Box Tuning
In scenarios where gradient access is unavailable, preliminary efforts have focused on federated fine-tuning black-box FMs (Lin et al., 2023; Sun et al., 2024a; Lu et al., 2023b; Rui et al., 2024) utilizing ZOO. However, ZOO’s noticeably slower convergence rates, especially in high-dimensional contexts compared to gradient-based methods (Golovin et al., 2020), indicate an important direction for further research. The impact of these slower convergence rates on overall efficiency and computational load within FL, particularly concerning large-scale FMs, has not been adequately investigated and understood.
FL with AI-Generated Content
AI-Generated Content (AIGC) denotes content produced via advanced generative FMs (Wu et al., 2023a). The strong generative capability of FMs offers the advantage of rapidly automating the creation of inexhaustible synthetic data. This capability positions AIGC as a valuable supplementary data source for model training and evaluation in many tasks (Xu et al., 2024c). Despite some efforts (Zhang et al., 2023b), more potential opportunities and challenges for AIGC-aided FL have yet to be explored.
10. Conclusions
In this survey, we have meticulously surveyed the intersection of FM and FL. We identified core challenges in efficiency, adaptability, and trustworthiness and proposed a comprehensive taxonomy of techniques in response to these challenges. In addition, we discussed future directions and applications in this research field, hoping to attract more breakthroughs in future research.